### v4

# Depth-Breadth Synergy in RLVR: Unlocking LLM Reasoning Gains with Adaptive Exploration

Abstract
Reinforcement Learning with Verifiable Reward (RLVR) has emerged as a powerful paradigm for unlocking reasoning capabilities in large language models, yet its full potential is hindered by two under-explored dimensions: Depth—the hardest problem a model can sample; Breadth—the number of instances consumed in a single iteration. We dissect the popular GRPO algorithm and reveal a systematic bias: the cumulative-advantage disproportionately weights samples with medium accuracy, while down-weighting the low-accuracy instances that are crucial for pushing reasoning boundaries. To rectify the depth neglect, we introduce Difficulty Adaptive Rollout Sampling (DARS), which re-weights hard problems through targeted multi-stage rollouts, thereby increasing the number of positive rollouts for hard problems. Empirically, naively enlarging rollout size only accelerates convergence and even hurts Pass@K. Our DARS, in contrast, delivers consistent Pass@K gains without extra inference cost at convergence. Just as we adaptively expanded the depth of exploration, we now ask whether aggressively scaling the breadth of training data can further amplify reasoning gains. To this end, we intensely scale batch size and replace PPO’s mini-batch iterations with full-batch updates over multiple epochs. Increasing breadth significantly enhances Pass@1 performance. Large-breadth training sustains high token-level entropy, indicating continued exploration and reduced gradient noise. We further present DARS-Breadth, which augments DARS with large breadth, and demonstrate simultaneous gains in Pass@K and Pass@1. The results confirm that breadth and adaptive exploration across depth operate as orthogonal dimensions in RLVR, which are key to unleashing the reasoning power of RLVR.
Figure 1: Training dynamics of Pass@1 and Pass@K performance. We show that our DARS significantly improves Pass@K performance and is complementary to breadth scaling to further improve Pass@1 performance.
1 Introduction
The emergence of reasoning-centric Large Language Models (LLMs) exemplified by OpenAI-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025), and Kimi-1.5 (Team et al., 2025), has pushed the frontier of LLM capability, especially for demanding tasks in complex reasoning such as mathematics and programming. Unlike conventional instruction tuning that relies on human-labeled data or RLHF pipelines that demand an auxiliary, well-trained reward model (Ouyang et al., 2022a; Achiam et al., 2023; Grattafiori et al., 2024), this leap is driven by large-scale Reinforcement Learning with Verifiable Rewards (RLVR; Guo et al. 2025; Zeng et al. 2025) for which correctness can be automatically and deterministically checked. The rewards of RLVR are granted solely when a model’s output matches the ground-truth mathematical answer or passes all unit tests for code, allowing scalable verification without manual labeling. RLVR is now regarded as a promising path toward self-evolving LLMs, potentially bringing us closer to more powerful intelligence.
However, existing RLVR frameworks inadequately address the interplay between exploration depth (difficulty scaling) and breadth (iteration instance quantity scaling), which leads to insufficient performance gain for both Pass@1 and Pass@K. In this paper, we conduct a systematic analysis of two under-exploited dimensions in RLVR: Depth and Breadth.
For the dimension of depth, our investigation reveals that existing methods of GRPO (Shao et al., 2024) and its variants (Yu et al., 2025; Liu et al., 2025b), while adept at estimating the advantage of a single rollout, are undermined by a distorted cumulative advantage at the group level. This distortion disproportionately allocates attention to instances of medium difficulty, neglecting high-difficulty instances indispensable for complex reasoning, as illustrated in Figure 3. This bias fundamentally limits depth, the hardest problems a model can learn to solve, and constrains Pass@K performance. To counteract this depth neglect, we propose Difficulty-Adaptive Rollout Sampling (DARS). DARS performs a lightweight first-stage rollout to estimate per-problem accuracies, then allocates additional compute via targeted multi-stage rollouts to low-accuracy problems. By expanded sampling on hard problems, DARS re-weights the cumulative advantage, making it easier for LLMs to learn ‘deep’ samples and improving Pass@K performance.
We further identify breadth as the instance quantity consumed in a single iteration. We observe that breadth has a significant impact on the LLM’s performance and continuous exploration capability, as shown in Figure 5. We significantly increase the training batch size and replace PPO-minibatch updates with full-batch updates for multiple PPO-epochs. This seemingly simple change dramatically improves Pass@1 and sustains high token-level entropy throughout training, suggesting that breadth acts as implicit entropy regularization that delays premature convergence. Importantly, the gains from breadth are complementary to those from depth: we present DARS-Breadth that combines our DARS with large-breadth training, producing simultaneous boosts in both Pass@K and Pass@1.
Our contributions can be summarized as follows:
  • We conduct a systematic analysis on depth and breadth in RLVR, and uncover the depth bias in GRPO: cumulative advantage silently underweights low-accuracy, high-difficulty samples, capping Pass@K performance.
  • We introduce DARS, which reallocates compute from medium difficulty problems to the hardest problems via multi-stage rollout sampling. DARS re-weights the cumulative advantage distribution and quantitatively expands the sparse reward signals for difficult problems. In practice, DARS significantly improves Pass@K over multiple benchmarks.
  • We further illustrate that large breadth in RLVR training matters for the Pass@1 performance. Moreover, by combining DARS with large breadth training, we reveal the complementarity of Depth and Breadth in RLVR and acquire simultaneous boosts in both Pass@K and Pass@1 performance.
2 Understanding RLVR from Depth and Breadth
In this section, we identify the two dimensions of Depth and Breadth and examine their impact on the Pass@1 and Pass@K metrics during the RLVR process. For training details and specific experimental settings, please refer to Section 4.1.
2.1 Depth: The Hardest Problem Sampled in RLVR
We first identify Depth as the hardest problem that can be correctly answered in the RLVR training process. In the GRPO training process, groups whose entire rollouts yield incorrect answers suffer from gradient vanishing. Hence, sampling high-difficulty questions with correct reasoning paths is crucial for LLM training. In this subsection, we first show that merely increasing rollout size does not consistently yield significant gains in Pass@K performance, and sometimes can even be harmful. We then quantify GRPO’s cumulative advantage and highlight its under-weighting of high-difficulty samples.
Figure 2: Training dynamics of Pass@1 and Pass@K performance of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b with different rollout size.
Naive Scaling of Rollout Size Benefits Pass@1, But Not Necessarily Pass@K.
We present the training dynamics of Pass@1 and Pass@K performance during the RLVR training process in Figure 2. Enlarging the rollout size allows the sampling of correct solutions to hard problems during training. We originally assumed this would benefit Pass@K performance; however, experimental results show that this is not always the case. We find that Qwen2.5-Math-7b can significantly benefit from an increased rollout size, whereas for Qwen2.5-Math-1.5b, naively scaling rollout size can even harm Pass@K performance.
Figure 3: Statistical results of cumulative advantage. Group relative advantage calculation methods underestimate high-difficulty problems. $n$ denotes group size.
Cumulative Advantage Bias in GRPO Variants hinders the improvement of Pass@K.
In the GRPO framework, the advantage estimation is derived by normalizing binary rewards:
$\hat{A}_{i}^{std}=\frac{r_{i}-u}{\sigma},~~~~~\hat{A}_{i}^{nostd}=r_{i}-u,$ (1)
where $r_{i}$ is the binary reward of $i_{th}$ rollout, $u$ is the mean value of the group rewards $u=\text{mean}(\{R_{i}\}_{i=1}^{G})$ and $\sigma$ is the standard deviation of the group rewards $\sigma=\text{std}(\{R_{i}\}_{i=1}^{G})$. In the case of binary rewards, $u$ also represents the accuracy of LLM rollouts. Dr. GRPO  (Liu et al., 2025b) removes the standard-deviation term from the advantage computation to eliminate question-level difficulty bias, and demonstrates its superiority through extensive experiments. Consequently, the experiments reported in this study were conducted primarily though the Dr. GRPO methodology. We show more results of std-based advantage in Appendix C.1.
For a group $G$ with rollout size $N$, we define the cumulative advantage of a group as the sum of the absolute values of sample advantages: $\mathcal{A}_{\text{group}}\;=\;\sum_{i=1}^{G}\,|\hat{A}_{i}|$. The cumulative advantage reflects how much the algorithm weights each sample. Specifically, for Dr. GRPO,
$\mathcal{A}_{\text{group}}=2Nu(1-u),$ (2)
The cumulative advantage functional curve is plotted in Figure 3. As shown in the figure, group-based advantage computation funnels its weight toward problems of medium difficulty while largely overlooking those that are highly difficult. This bias limits the Pass@K performance of RLVR.
2.2 Breadth: Iteration Instance Quantity in RLVR
We identify Breadth as the quantity of instances used in each iteration of the RLVR process. In this subsection, we discuss the Pass@1 performance benefits of simply increasing the batch size for the RLVR process. In section 4.3, we further discussed the impact of replacing the PPO mini batch with the full batch update for multiple PPO epochs.
Breadth Matters for Pass@1 Performance.
Most studies (Liu et al., 2025b; a; Yan et al., 2025; Fu et al., 2025) conventionally set the batch size to 128. In this subsection, we drastically increase the batch size to 3072 and plot the training dynamics of Pass@1 and Pass@32 performance in Figure 4. Naively increasing the batch size brings a Pass@1 improvement for all models, yet it harms the Pass@128 performance of Qwen2.5-Math-1.5b. We consider that increasing the quantity of instances used in each iteration makes the gradient direction more accurate and reduces the impact of noise, thereby improving Pass@1 performance.
Figure 4: Training dynamics of Pass@1 and Pass@K performance of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b with different batch size.
Figure 5: Training dynamics of Pass@1 performance and token entropy for Qwen2.5-Math-1.5b and Qwen2.5-Math-7b.
Breadth Sustains Entropy for Model Exploration.
We further analyze the relationship between Pass@1 and token entropy during the training process, as shown in Figure 5. Higher token entropy signifies stronger exploration capabilities in LLMs. Increasing training breadth allows an LLM to achieve higher entropy at the same Pass@1 accuracy. We consider that large breadth serves as an entropy regularization mechanism, preventing premature convergence and boosting Pass@1 while sustaining entropy.
3 Methodology
In Section 2, we analyze the bias inherent in group-based advantage computation. To solve this issue, we introduce Difficulty Adaptive Rollout Sampling (DARS), which rebalances the cumulative advantage via multi-stage sampling. By further synergizing the depth and breadth training dimensions, we propose DARS-B, which improves both Pass@1 and Pass@K.
3.1 Difficulty Adaptive Rollout Sampling (DARS)
As shown in Figure 6, given a data batch $\mathcal{B}=\{q_{j}\}_{j=1}^{M}$ of reasoning questions, DARS operates in two phases: (i) pre-rollout difficulty estimation that assigns to each question $q_{j}$ a scalar difficulty score $x_{j}\in[0,1]$; and (ii) multi-stage rollout re-balancing that dynamically decides how many additional trajectories $\Delta n_{j}$ shall be allocated to $q_{j}$ so that the cumulative advantage for low-accuracy problems is up-weighted. To simplify the subsequent formula representation, we define
$\mathcal{S}(\hat{a}_{j})=2\hat{a}_{j}(1-\hat{a}_{j}).$ (3)
Phase 1: Pre-Rollout Difficulty Estimation. For every $q_{j}$, we draw a light first-stage rollout consisting of $N^{pre}$ independent trajectories $\{\tau_{j}^{(i)}\}$. Let the per-trajectory reward be binary, $r_{j}^{(i)}\in\{0,1\}$. We define the empirical accuracy
$\hat{a}_{j}=\frac{1}{N^{pre}}\sum_{i=1}^{N^{pre}}r_{j}^{(i)}.$ (4)
The difficulty score is then set to the complementary accuracy $x_{j}=1-\hat{a}_{j}$, so that $x_{j}\approx 1$ for the hardest problems and $x_{j}\approx 0$ for the easiest ones.
Phase 2: Multi-Stage Rollout Re-Balancing. Let $\mathcal{A}_{\text{group}}^{N}(u)$ denote the cumulative advantage under GRPO for a group whose average accuracy is $u$ with rollout size $N$. We aim to reallocate $\Delta N$ additional trajectories across the mini-batch so that the effective cumulative advantage for each question becomes an increasing function of its difficulty. To control the computing cost, we cap the rollout sampling upper limit at $N^{\mathrm{max}}$. To this end, we design two rebalancing schedules.
Schedule 1: Equal-Treatment (ET). For every question $q_{j}$ we enforce the rebalanced cumulative advantage as:
$\mathcal{A}_{\text{group}}^{ET}(q_{j})=\mathcal{A}_{\text{group}}^{N^{pre}}(0.5).$ (5)
We raise the cumulative advantage of all difficulty problems ($\hat{a}_{j}<0.5$) to the level achieved by a medium-difficulty problem with accuracy $\hat{a}_{j}=0.5$. The required extra trajectories are
$\Delta n_{j}^{\text{ET}}=\mathrm{min}(\left\lceil\frac{\mathcal{A}_{\text{group}}^{N^{pre}}(0.5)-\mathcal{A}_{\text{group}}^{N^{pre}}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}\right\rceil,N^{\mathrm{max}}-N^{pre}).$ (6)
Schedule 2: Hardness-Weighted (HW). We now impose a monotonically increasing re-weighting that allocates more rollouts to lower-accuracy problems:
$\mathcal{A}_{\text{group}}^{HW}(q_{j})=2(1-\hat{a}_{j})\mathcal{A}_{\text{group}}^{N^{pre}}(0.5).$ (7)
This yields
$\Delta n_{j}^{\text{HW}}=\mathrm{min}(\left\lceil\frac{2x_{j}\cdot\mathcal{A}_{\text{group}}^{N^{pre}}(0.5)-\mathcal{A}_{\text{group}}^{N^{pre}}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}\right\rceil,N^{\mathrm{max}}-N^{pre}).$ (8)
Figure 6: The overall training framework of our Difficulty Adaptive Rollout Sampling (DARS) with breadth scaling. Our DARS consists of 2 phases: 1) a pre-rollout stage to evaluate the difficulty of the given question, and 2) a re-balancing rollout stage to adjust the cumulative advantage. For breadth scaling, we replace ppo minibatch as full batch with multiple ppo epochs.
3.2 Depth Synergy with Breadth Scaling
Our analysis in Section 2.2 empirically confirms the substantial Pass@1 improvements from large-breadth training. While DARS primarily optimizes training depth via multi-stage rollout rebalance, its dynamic batch-size adjustments preclude standard PPO-style mini-batch updates. To resolve this architectural constraint while leveraging breadth benefits, we replace PPO’s mini-batch updates with full-batch gradient descent across multiple PPO epochs, as illustrated in Figure 6. This modification ensures compatibility with DARS’s dynamic allocation while maximizing effective training breadth per optimization step. We term this integrated approach DARS-Breadth, unifying depth-adaptive sampling with breadth maximization.
Full-batch training offers two key advantages: (1) elimination of mini-batch gradient noise, and (2) sustained token-level exploration, acting as implicit regularization against premature convergence. The resulting framework demonstrates complementary gains—DARS improves Pass@K through depth optimization, while large-breadth training enhances Pass@1, highlighting their synergistic roles in RLVR optimization.
3.3 Training Target
We adopt the clipped objective of GRPO without the KL penalty term. Following Dr. GRPO, we likewise remove the response length handling from the GRPO target. Specifically, for a problem $q$ sampled in training data $\mathcal{D}$, the training target is formalized as:
$\mathcal{J}(\theta)=$ $\mathbb{E}_{(q\sim\mathcal{D},\{o_{i}\}_{i=1}^{\mathcal{G}}\sim\pi_{\theta_{\text{old}}}(q)}$ (9)
$\Bigg[\frac{1}{G}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\Bigg(\min\Big(r_{i,t}(\theta)\hat{A}_{i,t},\ \text{clip}\Big(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\Big)\hat{A}_{i,t}\Big))\Bigg)\Bigg],$
where
$r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}.$ (10)
The token advantage $\hat{A}_{i,t}$ is computed using Equation 1.
4 Experiments
4.1 Setup
Evaluation and Training.
We evaluate the RLVR process using 5 widely used mathematical reasoning benchmarks: MATH-500 (Lightman et al., 2023), OlympiadBench (He et al., 2024), MinvervaMath (Lewkowycz et al., 2022), AIME24, and AMC23. We combine all of the evaluation benchmarks to report Pass@1 (Avg@128) and Pass@K performance. The training data used in this work is OpenR1-45K, which is a subset of OpenR1-Math-220k (Hugging Face, 2025). More details are shown in Appendix B.1.
Baseline and Methods:
We compare the following methods: (1) Other Post-Training Models: The Qwen2.5-Math-Instruct (Yang et al., 2024) series, SimpleRL-Zero-7B Zeng et al. (2025) model, and Oat-Zero (Liu et al., 2025b) series. (2) RLVR-baseline: Dr. GRPO with rollout size 8 and batch size 128. (3) Depth-Naive: Simply increasing the rollout size to 32. (4) Breadth-Naive: Simply increasing the batch size to 3072. (5) DARS-ET/HW: Our algorithm introduced in Section 3.1 with Equal-Treat/Hardness-Weighted schedule, using batch size 128 and $N^{\max}=32$. (6) DARS-ET/HW-Breadth: Our Depth-and-Breadth synergy algorithm introduced in Section 3.2, using batch size 3072 and $N^{\max}=32$. For all methods, the number of PPO mini-steps is uniformly set to 2.
Evaluation Protocol:
For all baseline methods, we select the checkpoint with the best Pass@1 performance for reporting. For our DARS method, we selected the checkpoint that achieved the best Pass@128 performance after surpassing the baseline Pass@1 performance. Table 1 summarizes the Avg@128 performance on each benchmark, the overall Pass@1 across all test data, and the Pass@128 performance.
Table 1: Overall performance of Pass@1 (Avg@128) and Pass@128 of Qwen2.5-Math series.
Base Model + Method AIME24 MATH500 Olympiad AMC Minerva Avg@128 Pass@128
Qwen2.5-Math-1.5B as the Base Model
Qwen2.5-Math-1.5B-Base 4.0 35.1 16.2 20.8 9.5 21.1 77.9
Qwen2.5-Math-1.5B-Ins 10.2 67.9 34.7 42.6 24.4 43.5 79.9
Oat-Zero-1.5B 16.4 73.0 35.5 47.4 26.8 46.3 80.3
Our RLVR Replication 14.7 75.9 39.4 47.5 31.2 49.6 79.6
Depth-Naive 16.5 76.2 39.9 47.9 30.9 50.1 79.9
Breadth-Naive 18.5 77.6 41.7 49.8 31.9 51.5 79.2
DARS-1.5B-ET 15.8 76.0 40.9 47.2 30.0 50.0 81.2
DARS-1.5B-ET-Breadth 18.6 79.4 42.9 50.6 31.7 52.5 80.8
DARS-1.5B-HW 14.7 76.4 40.0 48.4 30.8 50.0 82.1
DARS-1.5B-HW-Breadth 19.3 79.0 42.7 51.9 31.6 52.4 82.2
Qwen2.5-Math-7B as the Base Model
Qwen2.5-Math-7B-Base 11.6 52.3 19.7 35.2 15.3 30.1 82.1
Qwen2.5-Math-7B-Ins 12.9 81.5 39.9 47.0 34.1 52.0 82.3
SimpleRL-Zero-7B 23.3 72.8 36.1 52.8 26.8 46.9 82.5
Oat-Zero-7B 31.3 79.2 42.5 59.4 33.7 53.4 79.7
Our RLVR Replication 26.8 82.2 44.3 57.2 35.7 55.3 81.4
Depth-Naive 28.0 83.8 46.4 59.0 37.3 57.0 80.3
Breadth-Naive 30.5 83.7 47.3 61.4 37.7 57.7 79.2
DARS-7B-ET 26.9 83.2 46.6 57.3 38.5 57.0 81.7
DARS-7B-ET-Breadth 33.3 83.8 47.8 61.3 38.4 58.1 82.1
DARS-7B-HW 30.1 83.5 47.1 59.4 37.2 57.3 83.5
DARS-7B-HW-Breadth 33.0 84.5 48.4 63.0 36.9 58.4 83.4
4.2 Main Results
Breadth scaling delivers a clear and consistent boost to Pass@1. Across every model scale and every benchmark, Breadth-Naive outperforms both the GRPO baseline and Depth-Naive, lifting average Pass@1 (Avg@128) by 1.9–3.7 points on AIME24, MATH500, and Olympiad tasks. This advantage is not merely additive: when breadth is combined with depth through DARS-Breadth, the margin widens further. DARS-Breadth reliably beats both Breadth-Naive and the original DARS variants, confirming our central hypothesis—depth and breadth are complementary, not competing, resources. Their synergy is what unlocks the next tier of LLM reasoning gains.
The practical impact is twofold. First, DARS-Breadth secures the highest Pass@1, the metric that matters most for single-shot deployment. Second, it matches the best Pass@128 scores, demonstrating that the breadth-depth collaboration does not sacrifice the upper-bound capability revealed by heavy sampling. Finally, the choice of schedule matters: the HW schedule consistently yields superior Pass@K curves for both breadth and non-breadth training, while maintaining Pass@1 parity with the ET schedule, making it the preferred option across the board.
4.3 Training Dynamics and Ablation Study
In this subsection, we further show more training dynamics to illustrate properties of existing RLVR methods and the superiority of our DARS and DARS-B.
Pass@128 performance surpasses the base model, peaks quickly, and then declines.
We conduct RLVR experiments with rollout size 8/32 to compare our DARS (with $N^{\mathrm{max}}=32$), the training dynamics of Pass@128 performance during training is shown in Figure 7. Across all settings, Pass@128 surpasses the base model during training, but declines after peaking, indicating that over-training with RLVR harms Pass@128 performance. Notably, DARS (with $N^{\mathrm{max}}=32$) incurs substantially less inference cost than naively scaling the rollout size to n = 32. Despite this being an unfair comparison in terms of computational expenditure, our DARS not only attains the highest peak Pass@128 performance but also outperforms all other settings.
Figure 7: Training dynamics of Pass@128 performance with different training steps of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b.
Figure 8: Training dynamics of Pass@32/Pass@128 and Pass@1 performance with different training steps of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b.
Depth Training with DARS Improve Pass@K Performance and Training Efficiency.
Because the Pass@K (K=32/128) metric is hard to improve monotonically—it even starts to drop after prolonged training—while Pass@1 remains comparatively stable and rarely collapses, we seek to boost Pass@K without degrading Pass@1. Figure 8 plots Pass@128 against Pass@1 under a variety of experimental settings. It shows that, at any fixed Pass@1 level, our DARS method delivers a consistently higher Pass@128 than the other settings.
Table 2: Average rollout numbers per prompt.
Model Naive DARS-ET DARS-HW
Qwen2.5-Math-1.5B 32 15.2 ($\downarrow$52.5%) 23.9 ($\downarrow$25.3%)
Qwen2.5-Math-7B 32 12.8 ($\downarrow$60.0%) 20.1 ($\downarrow$37.2%)
It is worth noting that, unlike the naive approach of simply increasing the rollout size to 32, our DARS achieves significantly higher training efficiency by allocating more rollouts to the hard problems. As shown in Table 2, our DARS methods need far fewer rollouts than the Depth-Naive method while achieve better performance.
Depth-Breadth are Complementary in RLVR.
We further illustrate that Depth and Breadth are two complementary dimensions in RLVR. As shown in Figure 9, we present the Pass@1Pass@K training-dynamics curves for the Breadth, Depth, and the two-dimensional synergy approach DARS-Breadth. The farther the Pass@1Pass@K curve deviates outward, the more powerful the method. Our DARS-Breadth curves lie on the outermost envelope: it not only achieves the best Pass@1, but also simultaneously lifts Pass@K. This demonstrates the complementary roles of Depth and Breadth in RLVR training.
Figure 9: Complementary improve of Depth and Breadth Synergy for Pass@1 and Pass@K (K=128) performance.
Ablation Study on Base Model.
We further illustrate the effectiveness of our DARS method on Llama-3.1-8B. The results are shown in Table 3. Our DARS-ET-Breadth achieves both Pass@1 and Pass@128 performance compared to other baselines, which further illustrates the effectiveness of our method.
Table 3: Overall performance of Pass@1 (Avg@128) and Pass@128 performance of Llama-3.1-8B.
Base Model + Method AIME24 MATH500 Olympiad AMC Minerva Avg@128 Pass@128
Llama-3.1-8B 0.23 6.13 1.54 2.76 2.72 3.25 52.7
GRPO baseline 0.66 29.6 7.09 10.1 15.7 15.8 56.5
Depth-Naive 0.43 33.6 9.40 12.3 19.7 18.9 58.6
Breadth-Naive 0.79 34.4 9.34 12.2 19.0 19.0 61.1
DARS-Llama-ET-Breadth 1.46 39.4 12.0 13.2 20.1 22.0 67.2
DARS-Llama-HW-Breadth 1.11 39.0 12.0 13.3 19.8 21.8 68.7
Complete Pass@K Accuracy Curve.
We show the complete Pass@K curve for Llama-3.1-8B, Qwen2.5-Math-1.5B, and Qwen2.5-Math-7B in Figure 10. The 3 chosen models of DARS are: DARS-Llama-ET-Breadth, DARS-1.5B-HW-Breadth, and DARS-7B-HW-Breadth. DARS models demonstrate a breakthrough in the reasoning boundaries of the base model, especially on the LLama-3.1-8B model, where the improvement in Pass@k is particularly significant.
Figure 10: Complete Pass@K accuracy curve of base models and our DARS models.
5 Related Works
5.1 RL for LLM Reasoning
Reinforcement Learning (RL) is widely used in the post-training phase of LLMs. Previously, a reward model was employed to conduct preference comparisons among responses (Ouyang et al., 2022b). Direct Preference Optimization (Rafailov et al., 2023) is further proposed to simplify the training pipeline of RL for LLMs, which directly uses pair-wise preference data for model optimization. Recent studies have shown that RLVR can significantly improve the reasoning capabilities of LLMs in domains such as mathematics and programming. OpenAI’s o1 model (Jaech et al., 2024) represents a major milestone in LLM reasoning. Shortly thereafter, DeepSeek-R1 (Guo et al., 2025) demonstrated that RLVR can directly elicit a base model’s slow-thinking capability—a setting now termed zero-RL (Li et al., 2025). Subsequently, a series of Large Reasoning Models (LRM), including Kimi 1.5 (Team et al., 2025), Gemini-Think (DeepMind, 2024), and QwQ (Qwen, 2024) were successively introduced. These works have inspired a wave of follow-up research: Zeng et al. (2025) conducted extensive experimental exploration of zero-RL, while Luo et al. (2025) leveraged RLVR to further boost the performance of open-source models developed from DeepSeek-R1. Among these works, one widely adopted RLVR algorithm is GRPO (Shao et al., 2024), which builds upon PPO (Schulman et al., 2017) and attains substantial performance gains by sampling multiple responses to compute the group-relative advantage. GRPO has inspired a series of subsequent algorithms, including DAPO (Yu et al., 2025), VAPO (Yue et al., 2025b), and Dr. GRPO (Liu et al., 2025b). However, both GRPO and its variants tend to undervalue high-difficulty problems under cumulative advantage, which compromises the Pass@K performance of RLVR.
5.2 Analysis of RLVR
With the rapid advancement of RLVR and the proliferation of open-source LRMs, many studies have begun to analyze the RLVR pipeline and these open LRMs. Several studies (Liu et al., 2025a; Zhao et al., 2025; Shah et al., 2025) indicates that the self-reflect and self-critique behaviors observed after RLVR originates from the base model rather than the RL process. Dang et al. (2025) find that although the RLVR process benefits Pass@1, Pass@K may decline as training progresses. Subsequently, Yue et al. (2025a) through extensive experimental analysis, discovered that RLVR’s performance is significantly constrained by the base model; once training converges, it struggles to surpass the capability boundary of the base model. These studies have sparked widespread concern about the capability ceiling of RLVR, and consequently, the Pass@K metric has become a focal point for diagnosing and potentially transcending the intrinsic limits imposed by the base model (Liang et al., 2025). This paper analyzes and refines the RLVR pipeline from the dual perspectives of Pass@1 and Pass@K.
6 Conclusion
In this work, we identify and systematically analyze two critical yet under-explored dimensions in Reinforcement Learning with Verifiable Reward (RLVR): Depth (the hardest problems a model can learn) and Breadth (the number of instances processed per iteration). We reveal that existing GRPO-based methods suffer from a fundamental limitation—their cumulative advantage calculation disproportionately weights medium-difficulty problems while neglecting hard ones, creating a bottleneck for Pass@K performance. To address this issue, we propose Difficulty-Adaptive Rollout Sampling (DARS) to reallocate compute to the hard problems, boosting Pass@K with low extra inference cost. Furthermore, we demonstrate that large-breadth training acts as an orthogonal dimension for RLVR optimization. Scaling batch size and employing full-batch updates significantly boost Pass@1 performance while sustaining high token-level entropy. By unifying these dimensions in our DARS-Breadth framework, we achieve simultaneous gains in both Pass@1 and Pass@K, confirming the synergistic relationship between depth and breadth in RLVR.
Appendix
Appendix A Derivation of Additional Rollouts $\Delta n_{j}$
The cumulative advantage for a group with accuracy $\hat{a}_{j}$ and total rollout size $N_{j}={N^{pre}}+\Delta n_{j}$ is given by:
$\mathcal{A}_{\text{group}}(\hat{a}_{j},N_{j})=N_{j}\cdot\mathcal{S}(\hat{a}_{j}),$
where $\mathcal{S}(\hat{a}_{j})=2\hat{a}_{j}(1-\hat{a}_{j})$.
After the first-stage rollout of size $N^{pre}$}, the initial cumulative advantage is:
$\mathcal{A}_{\text{group}}^{N^{pre}}(\hat{a}_{j})={N^{pre}}\cdot\mathcal{S}(\hat{a}_{j}).$
Our goal is to determine the number of additional trajectories $\Delta n_{j}$ needed so that the final cumulative advantage $\mathcal{A}_{\text{group}}(\hat{a}_{j},N_{j})$ meets a target value $\mathcal{A}_{\text{group}}^{\text{target}}(\hat{a}_{j})$.
Equal-Treatment (ET) Schedule:
The target cumulative advantage is set to be constant for all questions with $\hat{a}_{j}<0.5$:
$\mathcal{A}_{\text{group}}^{\text{ET}}(\hat{a}_{j})=\mathcal{A}_{\text{group}}^{N^{pre}}(0.5)={N^{pre}}\cdot\mathcal{S}(0.5).$
We solve for $\Delta n_{j}^{\text{ET}}$:
$\mathcal{A}_{\text{group}}(\hat{a}_{j},N_{j})$ $=\mathcal{A}_{\text{group}}^{\text{ET}}(\hat{a}_{j})$
$({N^{pre}}+\Delta n_{j}^{\text{ET}})\cdot\mathcal{S}(\hat{a}_{j})$ $={N^{pre}}\cdot\mathcal{S}(0.5)$
$\Delta n_{j}^{\text{ET}}\cdot\mathcal{S}(\hat{a}_{j})$ $={N^{pre}}\cdot\mathcal{S}(0.5)-{N^{pre}}\cdot\mathcal{S}(\hat{a}_{j})$
$\Delta n_{j}^{\text{ET}}$ $=\frac{{N^{pre}}\cdot\mathcal{S}(0.5)-{N^{pre}}\cdot\mathcal{S}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}.$
$\Delta n_{j}^{\text{ET}}$ $=\frac{\mathcal{A}_{\text{group}}^{N^{pre}}(0.5)-\mathcal{A}_{\text{group}}^{N^{pre}}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}.$
The rollout size must be an integer, and we cap the total rollout sampling upper limit at $N^{max}$, so
$\Delta n_{j}^{\text{ET}}=\mathrm{min}(\left\lceil\frac{\mathcal{A}_{\text{group}}^{N^{pre}}(0.5)-\mathcal{A}_{\text{group}}^{N^{pre}}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}\right\rceil,N^{\mathrm{max}}-N^{pre}).$
Hardness-Weighted (HW) Schedule:
The target cumulative advantage increases with difficulty:
$\mathcal{A}_{\text{group}}^{\text{HW}}(\hat{a}_{j})=2(1-\hat{a}_{j})\cdot\mathcal{A}_{\text{group}}^{N}(0.5)=2x_{j}\cdot N^{pre}\cdot\mathcal{S}(0.5).$
We solve for $\Delta n_{j}^{\text{HW}}$:
$\mathcal{A}_{\text{group}}(\hat{a}_{j},N_{j})$ $=\mathcal{A}_{\text{group}}^{\text{HW}}(\hat{a}_{j})$
$(N^{pre}+\Delta n_{j}^{\text{HW}})\cdot\mathcal{S}(\hat{a}_{j})$ $=2x_{j}\cdot N^{pre}\cdot\mathcal{S}(0.5)$
$\Delta n_{j}^{\text{HW}}\cdot\mathcal{S}(\hat{a}_{j})$ $=2x_{j}\cdot N^{pre}\cdot\mathcal{S}(0.5)-N^{pre}\cdot\mathcal{S}(\hat{a}_{j})$
$\Delta n_{j}^{\text{HW}}$ $=\frac{2x_{j}\cdot N^{pre}\cdot\mathcal{S}(0.5)-N^{pre}\cdot\mathcal{S}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}.$
Again, using the baseline advantage notation $\mathcal{A}_{\text{group}}^{N^{pre}}(\hat{a}_{j})=N^{pre}\cdot\mathcal{S}(\hat{a}_{j})$, we obtain:
$\Delta n_{j}^{\text{HW}}=\mathrm{min}(\left\lceil\frac{2x_{j}\cdot\mathcal{A}_{\text{group}}^{N^{pre}}(0.5)-\mathcal{A}_{\text{group}}^{N^{pre}}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}\right\rceil,N^{\mathrm{max}}-N^{pre}).$
Both derivations include a ceiling function and are capped at $N^{\mathrm{max}}$ to control computational cost, as shown in Equations 6 and 8 in the paper.
Appendix B Training and Evaluation Details
B.1 Setting Details
[…]
Parameters and Metrics.
Currently, our experiments are conducted with Qwen2.5-Math series language models (Yang et al., 2024). We set the temperature to 1.0 for both the training and evaluation procedures. In this paper, we mainly use two metrics, Pass@1 and Pass@K. To acquire Pass@K results, we sample 128 candidate responses for each question during the evaluation process; the calculation of Pass@1 is derived from Avg@128. Both the training and evaluation processes are scored using Math-Verify. The learning rate is 1e-6 for depth training methods, and 5e-6 for large breadth training. We do not use the reference model and KL loss. For fair comparison, we uniformly set the PPO mini step to 2 for all experiments. By default, the maximum prompt length is 1024, and the maximum response length is 3072 for the Qwen2.5-Math series model.
Implementation Details.
Following LUFFY (Yan et al., 2025), we use the default subset and filter out generations that are longer than 8192 tokens and those that are verified wrong by Math-Verify  [^1] https://github.com/huggingface/Math-Verify, resulting in 45k question-solution pairs. For training Llama-3.1-8B, we use the train split of MATH dataset. Our training framework is derived from Verl (Sheng et al., 2024) pipeline, which is a flexible, high-performance reinforcement-learning framework built for training large language-model agents. With native PyTorch support and efficient distributed training, Verl lets researchers quickly prototype and scale RL algorithms like PPO on GPUs. All of our experiments are conducted on H200 GPUs. At present, the LLM of our experiment is the Qwen2.5-Math series.
Training Steps and Checkpoint Steps.
For non-breadth methods on Qwen2.5-Math-1.5B/7B, we set the checkpoint step as 100. For breadth methods on Qwen2.5-Math-1.5B/7B, we set the checkpoint step as 15. The specific training steps are determined according to the convergence of the model. The number of training steps for non-breadth training is set as 300 for Llama-3.1-8B, 600 for Qwen2.5-Math-1.5B, and 500 for Qwen2.5-Math-7B. The number of training steps for breadth training is set as 70. For breadth training, we set the total training steps as 105 for Qwen2.5-Math-1.5B, and 75 for Qwen2.5-Math-7B.
Appendix C More Experimental Results
C.1 Ablation Study on std-based Advantage Computation
As illustrated in Section 2, Dr. GRPO  (Liu et al., 2025b) removes the standard-deviation term from the advantage computation to eliminate question-level difficulty bias, and demonstrates its superiority through extensive experiments. Consequently, the experiments reported in this study were conducted primarily though the Dr. GRPO methodology. To further illustrate the effectiveness of DARS on std-based advantage computation, we conduct the experiment with HW schedule on Qwen2.5-Math-1.5B model, as shown in Figure 11
Figure 11: Comparison of our DARS on std-based advantage computation.
C.2 Depth and Breadth Synergy for Pass@1 and Pass@32
In Section 4.3, we show the training dynamics of Pass@128-Pass@1 for DARS and baseline methods. To further illustrate the effectiveness of DARS, we show the training dynamics of Pass@32-Pass@1 in Figure 12. Our DARS significantly improves the Pass@32 performance compared to other methods.
Figure 12: Complementary improve of Depth and Breadth Synergy for Pass@1 and Pass@K (K=32) performance.
C.3 Comparison of ET/HW Schedule in Breadth Scaling
In addition, compared with the ET schedule, DARS-HW-Breadth significantly improves the model’s Pass@128 performance as shown in Figure 13. We consider this performance gain is due to the HW schedule placing greater emphasis on difficult samples.
Figure 13: Comparison of ET and HW schedule in breadth training of Qwen2.5-Math series.
Appendix D Discussion and Future Work
In this section, we analyze how hyperparameters $N$ and $N^{\max}$ control the shape of the cumulative advantage curve, and how this shape may influence training behavior. We further discuss how dynamically adjusting these parameters could enable a smooth transition from Pass@K-oriented to Pass@1-oriented training.
D.1 Hyperparameter Control of Cumulative Advantage Shape
We show the Cumulative Advantage shape of ET/HW schedule with $N=8$ in Figure 14. By continuously reducing the size of $N_{\text{max}}$, the curve will contract accordingly. When $N_{\text{max}}=N$, it is equivalent to the vanilla method without performing DARS.
Figure 14: Control the shape of Cumulative Advantage by adjusting the $N_{max}$ hyperparameter of DARS.
D.2 Potential Pass@K to Pass@1 Training Transition
The dynamic control of $N^{\max}$ suggests an intriguing training strategy: starting with a large $N^{\max}$ value to maximize Pass@K performance through extensive exploration of hard problems, then gradually reducing $N^{\max}$ throughout training to transition toward Pass@1 optimization.
This approach mirrors curriculum learning principles, where the training difficulty is progressively adjusted. Initially, the model benefits from the expanded solution space and diverse reasoning patterns discovered through heavy sampling on hard problems (high $N^{\max}$). As training progresses and the model’s capability matures, reducing $N^{\max}$ focuses the training on refining the most promising solution strategies, ultimately improving single-shot performance.
Future work will explore optimal annealing schedules for $N^{\max}$ and investigate whether this transition strategy can simultaneously maximize both Pass@1 and Pass@K performance, potentially overcoming the current limitations of RLVR training.

### v5

@@
 Abstract
-Reinforcement Learning with Verifiable Reward (RLVR) has emerged as a powerful paradigm for unlocking reasoning capabilities in large language models, yet its full potential is hindered by two under-explored dimensions: Depth—the hardest problem a model can sample; Breadth—the number of instances consumed in a single iteration.
-We dissect the popular GRPO algorithm and reveal a systematic bias: the cumulative-advantage disproportionately weights samples with medium accuracy, while down-weighting the low-accuracy instances that are crucial for pushing reasoning boundaries.
-To rectify the depth neglect, we introduce Difficulty Adaptive Rollout Sampling (DARS), which re-weights hard problems through targeted multi-stage rollouts, thereby increasing the number of positive rollouts for hard problems.
-Empirically, naively enlarging rollout size only accelerates convergence and even hurts Pass@K.
-Our DARS, in contrast, delivers consistent Pass@K gains without extra inference cost at convergence.
-Just as we adaptively expanded the depth of exploration, we now ask whether aggressively scaling the breadth of training data can further amplify reasoning gains.
-To this end, we intensely scale batch size and replace PPO’s mini-batch iterations with full-batch updates over multiple epochs.
-Increasing breadth significantly enhances Pass@1 performance.
-Large-breadth training sustains high token-level entropy, indicating continued exploration and reduced gradient noise.
-We further present DARS-Breadth, which augments DARS with large breadth, and demonstrate simultaneous gains in Pass@K and Pass@1.
-The results confirm that breadth and adaptive exploration across depth operate as orthogonal dimensions in RLVR, which are key to unleashing the reasoning power of RLVR.
-Figure 1: Training dynamics of Pass@1 and Pass@K performance.
-We show that our DARS significantly improves Pass@K performance and is complementary to breadth scaling to further improve Pass@1 performance.
+Reinforcement Learning with Verifiable Reward (RLVR) is a powerful method for enhancing the reasoning abilities of Large Language Models, but its full potential is limited by a lack of exploration in two key areas: Depth (the difficulty of problems) and Breadth (the number of training instances).
+Our analysis of the popular GRPO algorithm reveals a bias that down-weights difficult, low-accuracy problems, which are crucial for improving reasoning skills.
+To address this, we introduce Difficulty Adaptive Rollout Sampling (DARS), a method that re-weights difficult problems by using targeted, multi-stage rollouts.
+DARS increases the number of rollout outcomes for these harder problems according to our proposed re-balancing schedules and leads to consistent gains in Pass@K.
+We discovered that increasing rollout size alone does not improve performance and may actually impair it.
+In contrast, scaling the batch size to increase breadth via full-batch updates significantly boosted Pass@1 metrics.
+This improvement stems from higher token-level entropy, ensuring robust exploration and minimized gradient noise.
+Finally, we present DARS-Breadth, a combined approach that uses DARS with a large breadth of training data.
+This method demonstrates simultaneous gains in both Pass@K and Pass@1, confirming that depth (adaptive exploration) and breadth (scaling the training data) are orthogonal and essential dimensions for unlocking the full power of RLVR.
+Machine Learning, ICML
+yangzhch6@gmail
+Project Repo: https://github.com/yangzhch6/DARS
 1 Introduction
 The emergence of reasoning-centric Large Language Models (LLMs) exemplified by OpenAI-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025), and Kimi-1.5 (Team et al., 2025), has pushed the frontier of LLM capability, especially for demanding tasks in complex reasoning such as mathematics and programming.
-Unlike conventional instruction tuning that relies on human-labeled data or RLHF pipelines that demand an auxiliary, well-trained reward model (Ouyang et al., 2022a; Achiam et al., 2023; Grattafiori et al., 2024), this leap is driven by large-scale Reinforcement Learning with Verifiable Rewards (RLVR; Guo et al.
+Unlike conventional instruction tuning that relies on human-labeled data or RLHF pipelines that demand an auxiliary, well-trained reward model (Ouyang et al., 2022b; Achiam et al., 2023; Grattafiori et al., 2024), this leap is driven by large-scale Reinforcement Learning with Verifiable Rewards (RLVR; Guo et al.
 2025; Zeng et al.
@@
 However, existing RLVR frameworks inadequately address the interplay between exploration depth (difficulty scaling) and breadth (iteration instance quantity scaling), which leads to insufficient performance gain for both Pass@1 and Pass@K.
-In this paper, we conduct a systematic analysis of two under-exploited dimensions in RLVR: Depth and Breadth.
+In this paper, we conduct a systematic analysis of these two under-exploited dimensions in RLVR: Depth and Breadth.
 For the dimension of depth, our investigation reveals that existing methods of GRPO (Shao et al., 2024) and its variants (Yu et al., 2025; Liu et al., 2025b), while adept at estimating the advantage of a single rollout, are undermined by a distorted cumulative advantage at the group level.
@@
 DARS performs a lightweight first-stage rollout to estimate per-problem accuracies, then allocates additional compute via targeted multi-stage rollouts to low-accuracy problems.
-By expanded sampling on hard problems, DARS re-weights the cumulative advantage, making it easier for LLMs to learn ‘deep’ samples and improving Pass@K performance.
+By expanding sampling on hard problems, DARS re-weights the cumulative advantage, making it easier for LLMs to learn ‘deep’ samples and improving Pass@K performance.
+Figure 1: Statistical results of cumulative advantage. Group relative advantage calculation methods underestimate high-difficulty problems. $n$ denotes group size.
 We further identify breadth as the instance quantity consumed in a single iteration.
@@
 Our contributions can be summarized as follows:
-  • We conduct a systematic analysis on depth and breadth in RLVR, and uncover the depth bias in GRPO: cumulative advantage silently underweights low-accuracy, high-difficulty samples, capping Pass@K performance.
+  • We conduct a systematic analysis on depth and breadth in RLVR, and uncover the depth bias in GRPO: cumulative advantage underweights low-accuracy, high-difficulty samples, capping Pass@K performance.
   • We introduce DARS, which reallocates compute from medium difficulty problems to the hardest problems via multi-stage rollout sampling.
@@
 2 Understanding RLVR from Depth and Breadth
-In this section, we identify the two dimensions of Depth and Breadth and examine their impact on the Pass@1 and Pass@K metrics during the RLVR process.
-For training details and specific experimental settings, please refer to Section 4.1.
+Figure 2: Training dynamics of Pass@1 and Pass@K performance of Qwen2.5-Math-1.5B and -7B with different rollout size.
 2.1 Depth: The Hardest Problem Sampled in RLVR
@@
 Hence, sampling high-difficulty questions with correct reasoning paths is crucial for LLM training.
-In this subsection, we first show that merely increasing rollout size does not consistently yield significant gains in Pass@K performance, and sometimes can even be harmful.
+We first show that merely increasing rollout size does not consistently yield significant gains in Pass@K performance, and sometimes can even be harmful.
 We then quantify GRPO’s cumulative advantage and highlight its under-weighting of high-difficulty samples.
-Figure 2: Training dynamics of Pass@1 and Pass@K performance of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b with different rollout size.
 Naive Scaling of Rollout Size Benefits Pass@1, But Not Necessarily Pass@K.
@@
 We find that Qwen2.5-Math-7B can significantly benefit from an increased rollout size, whereas for Qwen2.5-Math-1.5b, naively scaling rollout size can even harm Pass@K performance.
-Figure 3: Statistical results of cumulative advantage. Group relative advantage calculation methods underestimate high-difficulty problems. $n$ denotes group size.
-Cumulative Advantage Bias in GRPO Variants hinders the improvement of Pass@K.
-In the GRPO framework, the advantage estimation is derived by normalizing binary rewards:
+Cumulative Advantage Bias in GRPO Variants hinders the improvement of Pass@K. In the GRPO framework, the advantage estimation is derived by normalizing binary rewards:
 $\hat{A}_{i}^{std}=\frac{r_{i}-u}{\sigma},~~~~~\hat{A}_{i}^{nostd}=r_{i}-u,$ (1)
@@
 GRPO methodology.
-We show more results of std-based advantage in Appendix C.1.
+We show more results of std-based advantage in Appendix D.
+Figure 3: Training dynamics of Pass@1 and Pass@K performance of Qwen2.5-Math-1.5B and -7B with different batch size.
 For a group $G$ with rollout size $N$, we define the cumulative advantage of a group as the sum of the absolute values of sample advantages: $\mathcal{A}_{\text{group}}\;=\;\sum_{i=1}^{G}\,|\hat{A}_{i}|$.
@@
 GRPO,
-$\mathcal{A}_{\text{group}}=2Nu(1-u),$ (2)
+$\mathcal{A}_{\text{group}}=2Nu(1-u).$ (2)
 The cumulative advantage functional curve is plotted in Figure 1.
@@
 2.2 Breadth: Iteration Instance Quantity in RLVR
-We identify Breadth as the quantity of instances used in each iteration of the RLVR process.
-In this subsection, we discuss the Pass@1 performance benefits of simply increasing the batch size for the RLVR process.
-In section 4.3, we further discussed the impact of replacing the PPO mini batch with the full batch update for multiple PPO epochs.
+We define Breadth as the number of instances used per iteration of the RLVR process. We’ll show how increasing the batch size for the RLVR process improves the Pass@1 performance.
 Breadth Matters for Pass@1 Performance.
-Most studies (Liu et al., 2025b; a; Yan et al., 2025; Fu et al., 2025) conventionally set the batch size to 128.
+Most studies (Liu et al., 2025b, a; Yan et al., 2025; Fu et al., 2025) conventionally set the batch size to 128.
 In this subsection, we drastically increase the batch size to 3072 and plot the training dynamics of Pass@1 and Pass@32 performance in Figure 3.
@@
 We consider that increasing the quantity of instances used in each iteration makes the gradient direction more accurate and reduces the impact of noise, thereby improving Pass@1 performance.
-Figure 4: Training dynamics of Pass@1 and Pass@K performance of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b with different batch size.
 Figure 4: Training dynamics of Pass@1 performance and token entropy for Qwen2.5-Math-1.5B and Qwen2.5-Math-7B.
 Breadth Sustains Entropy for Model Exploration.
-We further analyze the relationship between Pass@1 and token entropy during the training process, as shown in Figure 5.
-Higher token entropy signifies stronger exploration capabilities in LLMs.
-Increasing training breadth allows an LLM to achieve higher entropy at the same Pass@1 accuracy.
-We consider that large breadth serves as an entropy regularization mechanism, preventing premature convergence and boosting Pass@1 while sustaining entropy.
+High token entropy in LLMs indicates strong exploration capabilities.
+Our analysis shows a relationship between Pass@1 and token entropy during training.
+As illustrated in Figure 4, increased training breadth enables LLMs to achieve higher entropy at a given Pass@1 accuracy.
+We believe a large training breadth acts as a form of entropy regularization, preventing premature convergence and boosting Pass@1 performance while maintaining high entropy.
+Figure 5: The overall training framework of our Difficulty Adaptive Rollout Sampling (DARS) with breadth scaling.
+Our DARS consists of 2 phases: 1) a pre-rollout stage to evaluate the difficulty of the given question, and 2) a re-balancing rollout stage to adjust the cumulative advantage.
+For breadth scaling, we replace PPO minibatch as full batch with multiple PPO epochs.
 3 Methodology
@@
 Schedule 2: Hardness-Weighted (HW). We now impose a monotonically increasing re-weighting that allocates more rollouts to lower-accuracy problems:
-$\mathcal{A}_{\text{group}}^{HW}(q_{j})=2(1-\hat{a}_{j})\mathcal{A}_{\text{group}}^{N^{pre}}(0.5).$ (7)
+$\mathcal{A}_{\text{group}}^{HW}(q_{j})=2(1-x_{j})\mathcal{A}_{\text{group}}^{N^{pre}}(0.5).$ (7)
 This yields
-$\Delta n_{j}^{\text{HW}}=\mathrm{min}(\left\lceil\frac{2x_{j}\cdot\mathcal{A}_{\text{group}}^{N^{pre}}(0.5)-\mathcal{A}_{\text{group}}^{N^{pre}}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}\right\rceil,N^{\mathrm{max}}-N^{pre}).$ (8)
-Figure 6: The overall training framework of our Difficulty Adaptive Rollout Sampling (DARS) with breadth scaling.
-Our DARS consists of 2 phases: 1) a pre-rollout stage to evaluate the difficulty of the given question, and 2) a re-balancing rollout stage to adjust the cumulative advantage.
-For breadth scaling, we replace ppo minibatch as full batch with multiple ppo epochs.
+$\Delta n_{j}^{\text{HW}}=$ $\mathrm{min}(\left\lceil\frac{2x_{j}\cdot\mathcal{A}_{\text{group}}^{N^{pre}}(0.5)-\mathcal{A}_{\text{group}}^{N^{pre}}(\hat{a}_{j})}{\mathcal{S}(\hat{a}_{j})}\right\rceil,$ (8)
+$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~N^{\mathrm{max}}-N^{pre}).$
 3.2 Depth Synergy with Breadth Scaling
@@
 The resulting framework demonstrates complementary gains—DARS improves Pass@K through depth optimization, while large-breadth training enhances Pass@1, highlighting their synergistic roles in RLVR optimization.
+Table 1: Overall performance of Pass@1 (Avg@128) and Pass@128 of Qwen2.5-Math series.
+Base Model + Method AIME24 MATH500 Olympiad AMC Minerva Avg@128 Pass@128
+Qwen2.5-Math-1.5B 4.0 35.1 16.2 20.8 9.5 21.1 77.9
+RLVR baseline 14.7 75.9 39.4 47.5 31.2 49.6 79.6
+Depth-Naive 16.5 76.2 39.9 47.9 30.9 50.1 79.9
+Breadth-Naive 18.5 77.6 41.7 49.8 31.9 51.5 79.2
+DARS-1.5B-ET 15.8 76.0 40.9 47.2 30.0 50.0 81.2
+DARS-1.5B-ET-Breadth 18.6 79.4 42.9 50.6 31.7 52.5 80.8
+DARS-1.5B-HW 17.7 76.4 40.0 48.4 30.8 50.0 82.1
+DARS-1.5B-HW-Breadth 19.3 79.0 42.7 51.9 31.6 52.4 82.2
+Qwen2.5-Math-7B 11.6 52.3 19.7 35.2 15.3 30.1 82.1
+RLVR baseline 26.8 82.2 44.3 57.2 35.7 55.3 81.4
+Depth-Naive 28.0 83.8 46.4 59.0 37.3 57.0 80.3
+Breadth-Naive 30.5 83.7 47.3 61.4 37.7 57.7 79.2
+DARS-7B-ET 26.9 83.2 46.6 57.3 38.5 57.0 81.7
+DARS-7B-ET-Breadth 33.3 83.8 47.8 61.3 38.4 58.1 82.1
+DARS-7B-HW 30.1 83.5 47.1 59.4 37.2 57.3 83.5
+DARS-7B-HW-Breadth 33.0 84.5 48.4 63.0 36.9 58.4 83.4
 3.3 Training Target
@@
 Specifically, for a problem $q$ sampled in training data $\mathcal{D}$, the training target is formalized as:
-$\mathcal{J}(\theta)=$ $\mathbb{E}_{(q\sim\mathcal{D},\{o_{i}\}_{i=1}^{\mathcal{G}}\sim\pi_{\theta_{\text{old}}}(q)}$ (9)
-$\Bigg[\frac{1}{G}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\Bigg(\min\Big(r_{i,t}(\theta)\hat{A}_{i,t},\ \text{clip}\Big(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\Big)\hat{A}_{i,t}\Big))\Bigg)\Bigg],$
+$\mathcal{J}(\theta)=\mathbb{E}_{(q\sim\mathcal{D},\{o_{i}\}_{i=1}^{\mathcal{G}}\sim\pi_{\theta_{\text{old}}}(q)}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\Bigg($ (9)
+$~~~~~\min\Big(r_{i,t}(\theta)\hat{A}_{i,t},\ \text{clip}\Big(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\Big)\hat{A}_{i,t}\Big)\Bigg)\Bigg],$
 where
@@
 The token advantage $\hat{A}_{i,t}$ is computed using Equation 1.
+Figure 6: Training dynamics of Pass@128 performance with different training steps of Qwen2.5-Math-1.5B and Qwen2.5-Math-7B.
 4 Experiments
 4.1 Setup
-Evaluation and Training.
-We evaluate the RLVR process using 5 widely used mathematical reasoning benchmarks: MATH-500 (Lightman et al., 2023), OlympiadBench (He et al., 2024), MinvervaMath (Lewkowycz et al., 2022), AIME24, and AMC23.
+Evaluation and Training: We evaluate the RLVR process using 5 widely used mathematical reasoning benchmarks: MATH-500 (Lightman et al., 2023), OlympiadBench (He et al., 2024), MinvervaMath (Lewkowycz et al., 2022), AIME24, and AMC23.
 We combine all of the evaluation benchmarks to report Pass@1 (Avg@128) and Pass@K performance.
 The training data used in this work is OpenR1-45K, which is a subset of OpenR1-Math-220k (Hugging Face, 2025).
-More details are shown in Appendix B.1.
-Baseline and Methods:
-We compare the following methods: (1) Other Post-Training Models: The Qwen2.5-Math-Instruct (Yang et al., 2024) series, SimpleRL-Zero-7B Zeng et al.
-(2025) model, and Oat-Zero (Liu et al., 2025b) series.
-(2) RLVR-baseline: Dr.
+Moreover, we adopt the same unbiased, low-variance estimator for Pass@K as used in prior works (Yue et al., 2025a; Chen et al., 2021):
+$\text{pass@}K=\mathbb{E}_{x_{i}\sim\mathcal{D}}\left[1-\frac{\binom{n-c_{i}}{k}}{\binom{n}{k}}\right$].
+Specifically, when $K=N$, the metric become: $\text{pass@}K=c_{1}\lor c_{2}\lor\cdots c_{128}$.
+Baseline and Methods: We compare with: (1) RLVR-baseline: Dr.
 GRPO with rollout size 8 and batch size 128.
@@
 For all methods, the number of PPO mini-steps is uniformly set to 2.
-Evaluation Protocol:
-For all baseline methods, we select the checkpoint with the best Pass@1 performance for reporting.
-For our DARS method, we selected the checkpoint that achieved the best Pass@128 performance after surpassing the baseline Pass@1 performance.
+Evaluation Protocol: For all baselines, we select the checkpoint with the best Pass@1 performance for reporting.
+For DARS, we selected the checkpoint that achieved the best Pass@128 performance after surpassing the baseline Pass@1 performance.
 Table 1 summarizes the Avg@128 performance on each benchmark, the overall Pass@1 across all test data, and the Pass@128 performance.
-Table 1: Overall performance of Pass@1 (Avg@128) and Pass@128 of Qwen2.5-Math series.
-Base Model + Method AIME24 MATH500 Olympiad AMC Minerva Avg@128 Pass@128
-Qwen2.5-Math-1.5B as the Base Model
-Qwen2.5-Math-1.5B-Base 4.0 35.1 16.2 20.8 9.5 21.1 77.9
-Qwen2.5-Math-1.5B-Ins 10.2 67.9 34.7 42.6 24.4 43.5 79.9
-Oat-Zero-1.5B 16.4 73.0 35.5 47.4 26.8 46.3 80.3
-Our RLVR Replication 14.7 75.9 39.4 47.5 31.2 49.6 79.6
-Depth-Naive 16.5 76.2 39.9 47.9 30.9 50.1 79.9
-Breadth-Naive 18.5 77.6 41.7 49.8 31.9 51.5 79.2
-DARS-1.5B-ET 15.8 76.0 40.9 47.2 30.0 50.0 81.2
-DARS-1.5B-ET-Breadth 18.6 79.4 42.9 50.6 31.7 52.5 80.8
-DARS-1.5B-HW 14.7 76.4 40.0 48.4 30.8 50.0 82.1
-DARS-1.5B-HW-Breadth 19.3 79.0 42.7 51.9 31.6 52.4 82.2
-Qwen2.5-Math-7B as the Base Model
-Qwen2.5-Math-7B-Base 11.6 52.3 19.7 35.2 15.3 30.1 82.1
-Qwen2.5-Math-7B-Ins 12.9 81.5 39.9 47.0 34.1 52.0 82.3
-SimpleRL-Zero-7B 23.3 72.8 36.1 52.8 26.8 46.9 82.5
-Oat-Zero-7B 31.3 79.2 42.5 59.4 33.7 53.4 79.7
-Our RLVR Replication 26.8 82.2 44.3 57.2 35.7 55.3 81.4
-Depth-Naive 28.0 83.8 46.4 59.0 37.3 57.0 80.3
-Breadth-Naive 30.5 83.7 47.3 61.4 37.7 57.7 79.2
-DARS-7B-ET 26.9 83.2 46.6 57.3 38.5 57.0 81.7
-DARS-7B-ET-Breadth 33.3 83.8 47.8 61.3 38.4 58.1 82.1
-DARS-7B-HW 30.1 83.5 47.1 59.4 37.2 57.3 83.5
-DARS-7B-HW-Breadth 33.0 84.5 48.4 63.0 36.9 58.4 83.4
+Figure 7: Training dynamics of Pass@32/Pass@128 and Pass@1 with different training steps of Qwen2.5-Math-1.5B and -7B.
 4.2 Main Results
@@
 Finally, the choice of schedule matters: the HW schedule consistently yields superior Pass@K curves for both breadth and non-breadth training, while maintaining Pass@1 parity with the ET schedule, making it the preferred option across the board.
-4.3 Training Dynamics and Ablation Study
-In this subsection, we further show more training dynamics to illustrate properties of existing RLVR methods and the superiority of our DARS and DARS-B.
-Pass@128 performance surpasses the base model, peaks quickly, and then declines.
-We conduct RLVR experiments with rollout size 8/32 to compare our DARS (with $N^{\mathrm{max}}=32$), the training dynamics of Pass@128 performance during training is shown in Figure 7.
-Across all settings, Pass@128 surpasses the base model during training, but declines after peaking, indicating that over-training with RLVR harms Pass@128 performance.
-Notably, DARS (with $N^{\mathrm{max}}=32$) incurs substantially less inference cost than naively scaling the rollout size to n = 32.
-Despite this being an unfair comparison in terms of computational expenditure, our DARS not only attains the highest peak Pass@128 performance but also outperforms all other settings.
-Figure 7: Training dynamics of Pass@128 performance with different training steps of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b.
-Figure 8: Training dynamics of Pass@32/Pass@128 and Pass@1 performance with different training steps of Qwen2.5-Math-1.5b and Qwen2.5-Math-7b.
-Depth Training with DARS Improve Pass@K Performance and Training Efficiency.
-Because the Pass@K (K=32/128) metric is hard to improve monotonically—it even starts to drop after prolonged training—while Pass@1 remains comparatively stable and rarely collapses, we seek to boost Pass@K without degrading Pass@1.
-Figure 8 plots Pass@128 against Pass@1 under a variety of experimental settings.
-It shows that, at any fixed Pass@1 level, our DARS method delivers a consistently higher Pass@128 than the other settings.
 Table 2: Average rollout numbers per prompt.
@@
 Qwen2.5-Math-7B 32 12.8 ($\downarrow$60.0%) 20.1 ($\downarrow$37.2%)
-It is worth noting that, unlike the naive approach of simply increasing the rollout size to 32, our DARS achieves significantly higher training efficiency by allocating more rollouts to the hard problems.
-As shown in Table 2, our DARS methods need far fewer rollouts than the Depth-Naive method while achieve better performance.
-Depth-Breadth are Complementary in RLVR.
-We further illustrate that Depth and Breadth are two complementary dimensions in RLVR.
-As shown in Figure 9, we present the Pass@1Pass@K training-dynamics curves for the Breadth, Depth, and the two-dimensional synergy approach DARS-Breadth.
-The farther the Pass@1Pass@K curve deviates outward, the more powerful the method.
-Our DARS-Breadth curves lie on the outermost envelope: it not only achieves the best Pass@1, but also simultaneously lifts Pass@K.
-This demonstrates the complementary roles of Depth and Breadth in RLVR training.
-Figure 9: Complementary improve of Depth and Breadth Synergy for Pass@1 and Pass@K (K=128) performance.
-Ablation Study on Base Model.
-We further illustrate the effectiveness of our DARS method on Llama-3.1-8B.
-The results are shown in Table 3.
-Our DARS-ET-Breadth achieves both Pass@1 and Pass@128 performance compared to other baselines, which further illustrates the effectiveness of our method.
+Figure 8: Depth and Breadth Synergy for Pass@1 and Pass@K (K=128) performance.
 Table 3: Overall performance of Pass@1 (Avg@128) and Pass@128 performance of Llama-3.1-8B.
@@
 DARS-Llama-HW-Breadth 1.11 39.0 12.0 13.3 19.8 21.8 68.7
+4.3 Training Dynamics and Ablation Study
+In this subsection, we further show more training dynamics to illustrate properties of existing RLVR methods and the superiority of our DARS and DARS-B.
+Pass@128 performance surpasses the base model, peaks quickly, and then declines.
+We conduct RLVR experiments with rollout size 8/32 to compare our DARS (with $N^{\mathrm{max}}=32$), the training dynamics of Pass@128 performance during training is shown in Figure 6.
+Across all settings, Pass@128 surpasses the base model during training, but declines after peaking, indicating that over-training with RLVR harms Pass@128 performance.
+Notably, DARS (with $N^{\mathrm{max}}=32$) incurs substantially less inference cost than naively scaling the rollout size to n = 32.
+Despite this being an unfair comparison in terms of computational expenditure, our DARS not only attains the highest peak Pass@128 performance but also outperforms all other settings.
+Figure 9: Complete Pass@K accuracy curve of base models and our DARS models.
+Depth Training with DARS Improve Pass@K Performance and Training Efficiency.
+Because the Pass@K (K=32/128) metric is hard to improve monotonically—it even starts to drop after prolonged training—while Pass@1 remains comparatively stable and rarely collapses, we seek to boost Pass@K without degrading Pass@1.
+Figure 7 plots Pass@128 against Pass@1 under a variety of experimental settings.
+It shows that, at any fixed Pass@1 level, our DARS method delivers a consistently higher Pass@128 than the other settings.
+It is worth noting that, unlike the naive approach of simply increasing the rollout size to 32, our DARS achieves significantly higher training efficiency by allocating more rollouts to the hard problems.
+As shown in Table 2, our DARS methods need far fewer rollouts than the Depth-Naive method while achieve better performance.
+Figure 10: Comparison of our DARS-HW and GRPO-Baseline on AIME 2024.
+We show the average response length and accuracy for Qwen2.5-Math-1.5B and Qwen2.5-Math-7B.
+After training with DARS, the models can achieve higher accuracy by utilizing longer thinking process.
+Depth-Breadth are Complementary in RLVR.
+We show that Depth and Breadth are two complementary dimensions in RLVR.
+As shown in Figure 8, we present the Pass@1Pass@K training-dynamics curves for the Breadth, Depth, and the two-dimensional synergy approach DARS-Breadth.
+The farther the Pass@1Pass@K curve deviates outward, the more powerful the method.
+Our DARS-Breadth curves lie on the outermost envelope: it not only achieves the best Pass@1, but also simultaneously lifts Pass@K.
+This demonstrates the complementary roles of Depth and Breadth.
+Ablation Study on Base Model.
+We further illustrate the effectiveness of DARS on Llama-3.1-8B.
+The results are shown in Table 3.
+Our DARS-ET-Breadth achieves both the highest Pass@1 and Pass@128 performance, which further illustrates the effectiveness of our method.
+DARS also consistently improves the performance of openPangu (Chen et al., 2025), as shown in Appendix E.
+DARS Elicits Longer Thought.
+We tracked the response length dynamics during the training of DARS and GRPO.
+As shown in Figure 10, on the challenging AIME 2024 benchmark, DARS consistently produces longer reasoning traces than the baseline and has higher accuracy.
+These results reveal that our DARS stimulates the model to perform deeper thinking to solve hard problems.
 Complete Pass@K Accuracy Curve.
@@
 DARS models demonstrate a breakthrough in the reasoning boundaries of the base model, especially on the LLama-3.1-8B model, where the improvement in Pass@k is particularly significant.
-Figure 10: Complete Pass@K accuracy curve of base models and our DARS models.
 5 Related Works
-5.1 RL for LLM Reasoning
-Reinforcement Learning (RL) is widely used in the post-training phase of LLMs.
-Previously, a reward model was employed to conduct preference comparisons among responses (Ouyang et al., 2022b).
-Direct Preference Optimization (Rafailov et al., 2023) is further proposed to simplify the training pipeline of RL for LLMs, which directly uses pair-wise preference data for model optimization.
-Recent studies have shown that RLVR can significantly improve the reasoning capabilities of LLMs in domains such as mathematics and programming.
-OpenAI’s o1 model (Jaech et al., 2024) represents a major milestone in LLM reasoning.
-Shortly thereafter, DeepSeek-R1 (Guo et al., 2025) demonstrated that RLVR can directly elicit a base model’s slow-thinking capability—a setting now termed zero-RL (Li et al., 2025).
-Subsequently, a series of Large Reasoning Models (LRM), including Kimi 1.5 (Team et al., 2025), Gemini-Think (DeepMind, 2024), and QwQ (Qwen, 2024) were successively introduced.
-These works have inspired a wave of follow-up research: Zeng et al.
-(2025) conducted extensive experimental exploration of zero-RL, while Luo et al.
-(2025) leveraged RLVR to further boost the performance of open-source models developed from DeepSeek-R1.
-Among these works, one widely adopted RLVR algorithm is GRPO (Shao et al., 2024), which builds upon PPO (Schulman et al., 2017) and attains substantial performance gains by sampling multiple responses to compute the group-relative advantage.
-GRPO has inspired a series of subsequent algorithms, including DAPO (Yu et al., 2025), VAPO (Yue et al., 2025b), and Dr.
+Reinforcement Learning (RL) is now standard in post-training LLMs.
+After early reward-model pipelines (Ouyang et al., 2022a), Direct Preference Optimization (Rafailov et al., 2023) streamlined training by exploiting pairwise preferences.
+RL with verifiable rewards (RLVR) has since pushed reasoning benchmarks in math and code, culminating in OpenAI’s o1 (Jaech et al., 2024) and the zero-RL breakthrough of DeepSeek-R1 (Guo et al., 2025).
+Follow-up Large Reasoning Models—Kimi 1.5 (Team et al., 2025), Gemini-Think (DeepMind, 2024), QwQ (Qwen, 2024), and studies like (Zeng et al., 2025; Luo et al., 2025) further validate RLVR.
+The leading algorithm, GRPO (Shao et al., 2024), extends PPO (Schulman et al., 2017) with group-relative advantages, inspiring DAPO (Yu et al., 2025), VAPO (Yue et al., 2025b), and Dr.
 GRPO (Liu et al., 2025b).
-However, both GRPO and its variants tend to undervalue high-difficulty problems under cumulative advantage, which compromises the Pass@K performance of RLVR.
-5.2 Analysis of RLVR
+Yet GRPO and its variants systematically undervalue hard problems, hurting Pass@K.
+More related works are shown in Appendix A.
+6 Conclusion
+In this work, we reveal that GRPO-based RLVR methods under-weight hard problems due to cumulative-advantage bias, capping Pass@K.
+Our DARS sampler cheaply re-allocates rollouts to these hard instances, while large-breadth training with full-batch updates raises Pass@1.
+The unified DARS-Breadth framework jointly lifts Pass@1 and Pass@K, proving depth and breadth are synergistic levers in RLVR.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here
+Appendix
+Appendix A More Related Works
 With the rapid advancement of RLVR and the proliferation of open-source LRMs, many studies have begun to analyze the RLVR pipeline and these open LRMs.
 Several studies (Liu et al., 2025a; Zhao et al., 2025; Shah et al., 2025) indicates that the self-reflect and self-critique behaviors observed after RLVR originates from the base model rather than the RL process.
-Dang et al.
-(2025) find that although the RLVR process benefits Pass@1, Pass@K may decline as training progresses.
-Subsequently, Yue et al.
-(2025a) through extensive experimental analysis, discovered that RLVR’s performance is significantly constrained by the base model; once training converges, it struggles to surpass the capability boundary of the base model.
+(Dang et al., 2025) find that although the RLVR process benefits Pass@1, Pass@K may decline as training progresses.
+Subsequently, (Yue et al., 2025a) through extensive experimental analysis, discovered that RLVR’s performance is significantly constrained by the base model; once training converges, it struggles to surpass the capability boundary of the base model.
 These studies have sparked widespread concern about the capability ceiling of RLVR, and consequently, the Pass@K metric has become a focal point for diagnosing and potentially transcending the intrinsic limits imposed by the base model (Liang et al., 2025).
 This paper analyzes and refines the RLVR pipeline from the dual perspectives of Pass@1 and Pass@K.
-6 Conclusion
-In this work, we identify and systematically analyze two critical yet under-explored dimensions in Reinforcement Learning with Verifiable Reward (RLVR): Depth (the hardest problems a model can learn) and Breadth (the number of instances processed per iteration).
-We reveal that existing GRPO-based methods suffer from a fundamental limitation—their cumulative advantage calculation disproportionately weights medium-difficulty problems while neglecting hard ones, creating a bottleneck for Pass@K performance.
-To address this issue, we propose Difficulty-Adaptive Rollout Sampling (DARS) to reallocate compute to the hard problems, boosting Pass@K with low extra inference cost.
-Furthermore, we demonstrate that large-breadth training acts as an orthogonal dimension for RLVR optimization.
-Scaling batch size and employing full-batch updates significantly boost Pass@1 performance while sustaining high token-level entropy.
-By unifying these dimensions in our DARS-Breadth framework, we achieve simultaneous gains in both Pass@1 and Pass@K, confirming the synergistic relationship between depth and breadth in RLVR.
-Appendix
-Appendix A Derivation of Additional Rollouts $\Delta n_{j}$
+Appendix B Derivation of Additional Rollouts $\Delta n_{j}$
 The cumulative advantage for a group with accuracy $\hat{a}_{j}$ and total rollout size $N_{j}={N^{pre}}+\Delta n_{j}$ is given by:
@@
 Both derivations include a ceiling function and are capped at $N^{\mathrm{max}}$ to control computational cost, as shown in Equations 6 and 8 in the paper.
-Appendix B Training and Evaluation Details
-B.1 Setting Details
+Appendix C Mathematical Derivation: Group Cumulative Advantage and GRPO Gradient Norm
+In this appendix, we provide a detailed mathematical derivation demonstrating the relationship between Group Cumulative Advantage and the gradient norm in GRPO.
+This derivation substantiates the claim that $\mathcal{A}_{\text{group}}$ serves as an effective indicator of the model’s attention to specific problems during training.
+C.1 GRPO Gradient Formulation
+The GRPO objective function and its gradient are given by (we can remove the clip operation for simplified analysis):
+$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\bigg],$ (11)
+$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\hat{A}_{i}\nabla_{\theta}\rho_{i,t}(\theta)\bigg]$
+$=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\hat{A}_{i}\frac{\nabla_{\theta}\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})}\bigg]$
+$=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\bigg].$ (12)
+C.2 Derivation of Gradient Norm Upper Bound
+We now derive the upper bound relationship between the gradient norm and Group Cumulative Advantage.
+The gradient norm of our adopted GRPO algorithm is shown as the following:
+$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|=\left\|\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right]\right\|.$ (13)
+Since the norm is a convex function and expectation is linear, by Jensen’s inequality:
+$\left\|\mathbb{E}[X]\right\|\leq\mathbb{E}[\|X\|]$ (14)
+Thus:
+$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq\mathbb{E}\left[\left\|\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\right].$ (15)
+Applying the triangle inequality to the inner summation:
+$\left\|\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}a_{i,t}\right\|\leq\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left\|a_{i,t}\right\|,$ (16)
+where:
+$a_{i,t}=\frac{1}{\sum_{i=1}^{G}|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t}).$ (17)
+Therefore:
+$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left\|\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\right].$ (18)
+Then we have,
+$\left\|\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|=|\rho_{i,t}(\theta)|\cdot|\hat{A}_{i}|\cdot\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|.$ (19)
+Thus:
+$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}|\rho_{i,t}(\theta)|\cdot|\hat{A}_{i}|\cdot\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\right].$ (20)
+We further take a boundedness assumption.
+in policy optimization, we assume the gradient log-probabilities $\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|$ are bounded.
+Furthermore, it’s vital to notice that the importance ratio $\rho_{i,t}(\theta)$ are also bounded through the clip operation in GRPO algorithm.
+Thus, there exists a constant $C>0$ such that:
+$|\rho_{i,t}(\theta)|\cdot\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\leq C.$ (21)
+Therefore:
+$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq C\cdot\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}|\hat{A}_{i}|\right].$ (22)
+Noting that $|\hat{A}_{i}|$ is independent of $t$ for fixed $i$:
+$\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}|\hat{A}_{i}|=\sum_{i=1}^{G}|o_{i}|\cdot|\hat{A}_{i}|.$ (23)
+Thus:
+$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq C\cdot\mathbb{E}\left[\frac{\sum_{i=1}^{G}|o_{i}|\cdot|\hat{A}_{i}|}{\sum_{i=1}^{G}|o_{i}|}\right].$ (24)
+In this paper, we define the Group Cumulative Advantage as:
+$\mathcal{A}_{\text{group}}=\sum_{i=1}^{G}|\hat{A}_{i}|.$ (25)
+By the weighted arithmetic mean inequality:
+$\frac{\sum_{i=1}^{G}|o_{i}|\cdot|\hat{A}_{i}|}{\sum_{i=1}^{G}|o_{i}|}\leq\frac{\sum_{i=1}^{G}|o_{i}|\cdot\sum_{i=1}^{G}|\hat{A}_{i}|}{\sum_{i=1}^{G}|o_{i}|}=\sum_{i=1}^{G}|\hat{A}_{i}|=\mathcal{A}_{\text{group}}.$ (26)
+Therefore:
+$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq C\cdot\mathbb{E}\left[\mathcal{A}_{\text{group}}\right].$ (27)
+C.3 Theoretical Implications
+This derivation establishes that Group Cumulative Advantage $\mathcal{A}_{\text{group}}$ provides an upper bound for the expected gradient norm in GRPO.
+Consequently: Larger $\mathcal{A}_{\text{group}}$ values indicate stronger gradient signals for the corresponding problem.
+The training process allocates more ”attention” to problems with higher $\mathcal{A}_{\text{group}}$ values during parameter updates.
+Therefore, we consider $\mathcal{A}_{\text{group}}$ serves as a mathematically grounded indicator of problem importance in GRPO training dynamics.
+This theoretical foundation validates the use of Group Cumulative Advantage as a meaningful metric for analyzing training behavior and problem prioritization in GRPO.
+Appendix D Training and Evaluation Details
 […]
@@
 With native PyTorch support and efficient distributed training, Verl lets researchers quickly prototype and scale RL algorithms like PPO on GPUs.
+Following Dr.
+GRPO (Liu et al., 2025b), we remove the KL loss and the length normalization in GRPO.
 All of our experiments are conducted on H200 GPUs.
 At present, the LLM of our experiment is the Qwen2.5-Math series.
+Figure 11: Comparison of our DARS on std-based advantage computation.
 Training Steps and Checkpoint Steps.
@@
 For breadth training, we set the total training steps as 105 for Qwen2.5-Math-1.5B, and 75 for Qwen2.5-Math-7B.
-Appendix C More Experimental Results
-C.1 Ablation Study on std-based Advantage Computation
+Appendix E More Experimental Results
+E.1 Ablation Study on std-based Advantage Computation
 As illustrated in Section 2, Dr.
@@
 To further illustrate the effectiveness of DARS on std-based advantage computation, we conduct the experiment with HW schedule on Qwen2.5-Math-1.5B model, as shown in Figure 11
-Figure 11: Comparison of our DARS on std-based advantage computation.
-C.2 Depth and Breadth Synergy for Pass@1 and Pass@32
-In Section 4.3, we show the training dynamics of Pass@128-Pass@1 for DARS and baseline methods.
+E.2 Depth and Breadth Synergy for Pass@1 and Pass@32
+In Section 10, we show the training dynamics of Pass@128-Pass@1 for DARS and baseline methods.
 To further illustrate the effectiveness of DARS, we show the training dynamics of Pass@32-Pass@1 in Figure 13.
 Our DARS significantly improves the Pass@32 performance compared to other methods.
+Figure 12: Comparison of ET and HW schedule in breadth training of Qwen2.5-Math series.
 Figure 13: Complementary improve of Depth and Breadth Synergy for Pass@1 and Pass@K (K=32) performance.
-C.3 Comparison of ET/HW Schedule in Breadth Scaling
+E.3 Comparison of ET/HW Schedule in Breadth Scaling
 In addition, compared with the ET schedule, DARS-HW-Breadth significantly improves the model’s Pass@128 performance as shown in Figure 12.
 We consider this performance gain is due to the HW schedule placing greater emphasis on difficult samples.
-Figure 13: Comparison of ET and HW schedule in breadth training of Qwen2.5-Math series.
-Appendix D Discussion and Future Work
+Figure 14: Heat map of Pass@128 of Qwen2.5-Math series in different temperatures.
+E.4 Impact of Temperature
+Some researches (Karan & Du, 2025; Qin et al., 2025; Ni et al., 2025) indicates that temperature matters in LLM reasoning.
+To further illustrate the performance improvement under different temperatures, we additionally add the above experiments.
+The results show that the improvement of our method is consistent over different temperatures.
+The results are shown in Figure 14.
+E.5 Consistent Improvement During RL Process
+To further show that our method consistently improve model performance, we calculated the mean of Pass@128 and Pass@32 for the last 3 checkpoints of each method, as shown in Table 4.
+Table 4: Average performance of Pass@1/32/128 for the last 3 checkpoints during training.
+Model Pass@1 Pass@32 Pass@128
+Qwen2.5-Math-1.5B as the Base Model
+GRPO-Baseline 48.8 74.7 80.8
+Depth-Naive 49.5 74.4 80.6
+Breadth-Naive 51.4 74.4 79.8
+DARS-HW 49.5 75.7 81.9
+DARS-HW-Breadth 52.4 76.4 82.1
+Qwen2.5-Math-7B as the Base Model
+GRPO-Baseline 55.1 76.9 81.8
+Depth-Naive 56.4 76.7 80.9
+Breadth-Naive 57.2 76.7 81.3
+DARS-HW 56.8 78.8 83.4
+DARS-HW-Breadth 58.3 79.1 83.7
+E.6 Performance of Other Model
+We further evaluate our method on Qwen2.5-7B-Instruct. Following (Liang et al., 2025), we change the training data to Math12k. The results are shown in Table 5.
+We further evaluate DARS with the openPangu-7B architecture on Huawei Ascend NPU, to investigate the generalization ability of DARS on different models and hardware architectures.
+The experimental results are reported in Table 6, and DARS improves the performance of the openPangu model consistently on all datasets.
+As the results show, our method still outperforms the baseline in both the Pass@1 and Pass@K metrics.
+Table 5: Overall performance of Pass@1 (Avg@128) and Pass@128 of Qwen2.5-7B-Instruct.
+Model AIME24 Math500 Olympiad AMC Minerva Avg@128 Pass@128
+Qwen2.5-7B-Instruct 11.9 72.3 37.1 42.2 31.9 47.2 80.3
+GRPO-baseline 14.2 74.8 37.6 43.4 33.4 48.6 78.8
+DARS-HW-Breadth 15.6 76.5 38.4 44.7 34.6 49.6 82.3
+Table 6: Overall performance of DARS with openPangu on Ascend NPU.
+Model AIME24 Math500 Olympiad AMC Minerva Avg@32 Pass@32
+openPangu-7B 36.99 82.71 48.82 51.61 37.0 57.54 79.92
+openPangu-7B + DARS 60.58 93.95 67.96 81.06 43.4 72.56 86.59
+E.7 Explanation of Significant Pass@k Gain for Llama-3
+The stronger gains for Llama-3 likely stem from its base model characteristics.
+Recent studies (Gandhi et al., 2025; Wang et al., 2025) suggest that base models like Llama-3 lack mid-training, causing many valid responses to appear only at higher sampling budgets.
+As shown in the table below, Llama-3-8B-Base exhibits a much larger performance gap when increasing (K) from 8 to 32/64 compared to Qwen models as shown in Table 7.
+This indicates that Llama-3 has more correct solutions “hidden” deeper in the sampling space.
+Our DARS method, by focusing on hard examples, effectively uncovers these “deep” correct samples, leading to more substantial Pass@K improvements for Llama-3.
+Table 7: Pass@K gap comparison of Qwen2.5-Math-1.5/7B and LLama3-8B-Base
+Model Pass@8 Pass@32 - Pass@8 Pass@64 - Pass@8
+Qwen2.5-Math-1.5B 56.5 12.4 17.1
+Qwen2.5-Math-7B 63.5 10.5 14.4
+Llama3-8B-Base 17.2 19.1 27.3
+E.8 Thinking Length Dynamics
+This section investigates how DARS influences the reasoning length of LLMs.
+We tracked the response length dynamics during the training of Qwen2.5-Math-1.5B and 7B models.
+Our experiments reveal two key observations: (1) The training process shows a clear trend of increasing generation length, as shown in Figure 15.
+(2) When evaluated on AIME 2024, models trained with DARS consistently produce longer reasoning traces than the baseline, as shown in Figure 16 (For steps in multiples of 100, we take the average of 128 sampling results.
+To increase the density of data points, we also sample 16 times for every intermediate 50-step interval and take the average).
+These results provide concrete evidence that our DARS method successfully stimulates the model to perform deeper and more thorough thinking.
+Figure 15: Training dynamics of response length for GRPO and DARS.
+Figure 16: Statistical results of response length on AIME 2024 for GRPO and DARS.
+Appendix F Discussion and Future Work
 In this section, we analyze how hyperparameters $N$ and $N^{\max}$ control the shape of the cumulative advantage curve, and how this shape may influence training behavior.
 We further discuss how dynamically adjusting these parameters could enable a smooth transition from Pass@K-oriented to Pass@1-oriented training.
-D.1 Hyperparameter Control of Cumulative Advantage Shape
+F.1 Hyperparameter Control of Cumulative Advantage Shape
 We show the Cumulative Advantage shape of ET/HW schedule with $N=8$ in Figure 17.
 By continuously reducing the size of $N_{\text{max}}$, the curve will contract accordingly.
-When $N_{\text{max}}=N$, it is equivalent to the vanilla method without performing DARS.
-Figure 14: Control the shape of Cumulative Advantage by adjusting the $N_{max}$ hyperparameter of DARS.
-D.2 Potential Pass@K to Pass@1 Training Transition
+When $N_{\text{max}}=N$, it is equivalent to the vanilla method without DARS.
+F.2 Potential Pass@K to Pass@1 Training Transition
 The dynamic control of $N^{\max}$ suggests an intriguing training strategy: starting with a large $N^{\max}$ value to maximize Pass@K performance through extensive exploration of hard problems, then gradually reducing $N^{\max}$ throughout training to transition toward Pass@1 optimization.
@@
 Future work will explore optimal annealing schedules for $N^{\max}$ and investigate whether this transition strategy can simultaneously maximize both Pass@1 and Pass@K performance, potentially overcoming the current limitations of RLVR training.
+Figure 17: Control the shape of Cumulative Advantage by adjusting the $N_{max}$ hyperparameter of DARS.

### v6

@@
 However, existing RLVR frameworks inadequately address the interplay between exploration depth (difficulty scaling) and breadth (iteration instance quantity scaling), which leads to insufficient performance gain for both Pass@1 and Pass@K.
-In this paper, we conduct a systematic analysis of these two under-exploited dimensions in RLVR: Depth and Breadth.
+In this paper, we conduct a systematic analysis of these two under-exploited dimensions in RLVR.
 For the dimension of depth, our investigation reveals that existing methods of GRPO (Shao et al., 2024) and its variants (Yu et al., 2025; Liu et al., 2025b), while adept at estimating the advantage of a single rollout, are undermined by a distorted cumulative advantage at the group level.
@@
   • We conduct a systematic analysis on depth and breadth in RLVR, and uncover the depth bias in GRPO: cumulative advantage underweights low-accuracy, high-difficulty samples, capping Pass@K performance.
-  • We introduce DARS, which reallocates compute from medium difficulty problems to the hardest problems via multi-stage rollout sampling.
+  • We introduce DARS, which reallocates compute to the hard problems via multi-stage rollout sampling.
 DARS re-weights the cumulative advantage distribution and quantitatively expands the sparse reward signals for difficult problems.
-In practice, DARS significantly improves Pass@K over multiple benchmarks.
-  • We further illustrate that large breadth in RLVR training matters for the Pass@1 performance.
+In practice, DARS significantly improves Pass@K performance.
+We elaborate on the impact of cumulative advantage on the gradient norm in Appendix B.2, and further discuss the relationship between DARS-HW and MaxRL (Tajwar et al., 2026) in Appendix C.
+  • We further illustrate that large breadth matters for the Pass@1 performance.
 Moreover, by combining DARS with large breadth training, we reveal the complementarity of Depth and Breadth in RLVR and acquire simultaneous boosts in both Pass@K and Pass@1 performance.
@@
 This paper analyzes and refines the RLVR pipeline from the dual perspectives of Pass@1 and Pass@K.
-Appendix B Derivation of Additional Rollouts $\Delta n_{j}$
+Appendix B Mathematical Derivations for DARS and GRPO Training Dynamics
+B.1 Derivation of Additional Rollouts $\Delta n_{j}$
 The cumulative advantage for a group with accuracy $\hat{a}_{j}$ and total rollout size $N_{j}={N^{pre}}+\Delta n_{j}$ is given by:
@@
 Both derivations include a ceiling function and are capped at $N^{\mathrm{max}}$ to control computational cost, as shown in Equations 6 and 8 in the paper.
-Appendix C Mathematical Derivation: Group Cumulative Advantage and GRPO Gradient Norm
-In this appendix, we provide a detailed mathematical derivation demonstrating the relationship between Group Cumulative Advantage and the gradient norm in GRPO.
-This derivation substantiates the claim that $\mathcal{A}_{\text{group}}$ serves as an effective indicator of the model’s attention to specific problems during training.
-C.1 GRPO Gradient Formulation
-The GRPO objective function and its gradient are given by (we can remove the clip operation for simplified analysis):
-$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\bigg],$ (11)
-$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\hat{A}_{i}\nabla_{\theta}\rho_{i,t}(\theta)\bigg]$
-$=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\hat{A}_{i}\frac{\nabla_{\theta}\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})}\bigg]$
-$=\mathbb{E}_{q\sim\mathcal{D},\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\bigg].$ (12)
-C.2 Derivation of Gradient Norm Upper Bound
-We now derive the upper bound relationship between the gradient norm and Group Cumulative Advantage.
-The gradient norm of our adopted GRPO algorithm is shown as the following:
-$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|=\left\|\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right]\right\|.$ (13)
-Since the norm is a convex function and expectation is linear, by Jensen’s inequality:
-$\left\|\mathbb{E}[X]\right\|\leq\mathbb{E}[\|X\|]$ (14)
-Thus:
-$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq\mathbb{E}\left[\left\|\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\right].$ (15)
-Applying the triangle inequality to the inner summation:
-$\left\|\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}a_{i,t}\right\|\leq\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left\|a_{i,t}\right\|,$ (16)
-where:
-$a_{i,t}=\frac{1}{\sum_{i=1}^{G}|o_{i}|}\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t}).$ (17)
-Therefore:
-$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left\|\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\right].$ (18)
-Then we have,
-$\left\|\rho_{i,t}(\theta)\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|=|\rho_{i,t}(\theta)|\cdot|\hat{A}_{i}|\cdot\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|.$ (19)
-Thus:
-$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}|\rho_{i,t}(\theta)|\cdot|\hat{A}_{i}|\cdot\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\right].$ (20)
-We further take a boundedness assumption.
-in policy optimization, we assume the gradient log-probabilities $\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|$ are bounded.
-Furthermore, it’s vital to notice that the importance ratio $\rho_{i,t}(\theta)$ are also bounded through the clip operation in GRPO algorithm.
-Thus, there exists a constant $C>0$ such that:
-$|\rho_{i,t}(\theta)|\cdot\left\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right\|\leq C.$ (21)
-Therefore:
-$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq C\cdot\mathbb{E}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}|\hat{A}_{i}|\right].$ (22)
-Noting that $|\hat{A}_{i}|$ is independent of $t$ for fixed $i$:
-$\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}|\hat{A}_{i}|=\sum_{i=1}^{G}|o_{i}|\cdot|\hat{A}_{i}|.$ (23)
-Thus:
-$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq C\cdot\mathbb{E}\left[\frac{\sum_{i=1}^{G}|o_{i}|\cdot|\hat{A}_{i}|}{\sum_{i=1}^{G}|o_{i}|}\right].$ (24)
-In this paper, we define the Group Cumulative Advantage as:
-$\mathcal{A}_{\text{group}}=\sum_{i=1}^{G}|\hat{A}_{i}|.$ (25)
-By the weighted arithmetic mean inequality:
-$\frac{\sum_{i=1}^{G}|o_{i}|\cdot|\hat{A}_{i}|}{\sum_{i=1}^{G}|o_{i}|}\leq\frac{\sum_{i=1}^{G}|o_{i}|\cdot\sum_{i=1}^{G}|\hat{A}_{i}|}{\sum_{i=1}^{G}|o_{i}|}=\sum_{i=1}^{G}|\hat{A}_{i}|=\mathcal{A}_{\text{group}}.$ (26)
-Therefore:
-$\left\|\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)\right\|\leq C\cdot\mathbb{E}\left[\mathcal{A}_{\text{group}}\right].$ (27)
-C.3 Theoretical Implications
-This derivation establishes that Group Cumulative Advantage $\mathcal{A}_{\text{group}}$ provides an upper bound for the expected gradient norm in GRPO.
-Consequently: Larger $\mathcal{A}_{\text{group}}$ values indicate stronger gradient signals for the corresponding problem.
-The training process allocates more ”attention” to problems with higher $\mathcal{A}_{\text{group}}$ values during parameter updates.
-Therefore, we consider $\mathcal{A}_{\text{group}}$ serves as a mathematically grounded indicator of problem importance in GRPO training dynamics.
-This theoretical foundation validates the use of Group Cumulative Advantage as a meaningful metric for analyzing training behavior and problem prioritization in GRPO.
+B.2 Connection of Expected Squared Gradient Norm and Group Cumulative Advantage
+In this subsection, we will prove that: in the setting of group-relative advantages, the expected squared gradient norm allocated to a problem is proportional to the group cumulative advantage $\mathcal{A}_{\text{group}}$.
+Proposition B.1 In a Reinforcement Learning with Verifiable Reward (RLVR) setting utilizing group-relative advantages, the expected squared gradient norm of a problem group is directly proportional to its Group Cumulative Advantage $\mathcal{A}_{\text{group}}$.
+Proof.
+By definition, for a specific problem $q$, the group gradient $\mathbf{g}_{\text{group}}$ before taking the dataset expectation is formulated at the trajectory level as:
+$\mathbf{g}_{\text{group}}=\frac{1}{G}\sum_{i=1}^{G}\hat{A}_{i}\nabla_{\theta}\log\pi_{\theta}(o_{i}|q),$ (11)
+where $o_{i}$ denotes the $i$-th sampled trajectory.
+Let $\mathbf{g}_{i}\triangleq\nabla_{\theta}\log\pi_{\theta}(o_{i}|q)$ denote the score function of the trajectory.
+To evaluate the optimization step size, we compute the expected squared $L_{2}$ norm of the group gradient:
+$\mathbb{E}\left[\|\mathbf{g}_{\text{group}}\|^{2}\right]=\frac{1}{G^{2}}\mathbb{E}\left[\sum_{i=1}^{G}\hat{A}_{i}^{2}\|\mathbf{g}_{i}\|^{2}+\sum_{i\neq j}\hat{A}_{i}\hat{A}_{j}\langle\mathbf{g}_{i},\mathbf{g}_{j}\rangle\right].$ (12)
+While the cross terms $\mathbb{E}[\hat{A}_{i}\hat{A}_{j}\langle\mathbf{g}_{i},\mathbf{g}_{j}\rangle]$ do not strictly vanish (since the advantages $\hat{A}_{i}$ and $\hat{A}_{j}$ are coupled through the group mean reward and correlated with the score functions), they are heavily dominated by the diagonal terms in the context of Large Language Models (LLMs).
+In the enormously high-dimensional parameter space typical of LLMs, the policy gradient estimator exhibits notoriously high variance.
+The individual trajectory gradients $\mathbf{g}_{i}$ are highly stochastic and point in nearly orthogonal directions.
+Consequently, the inner product of different trajectory gradients $\langle\mathbf{g}_{i},\mathbf{g}_{j}\rangle$ is exceedingly small compared to the squared norm of a single trajectory gradient $\|\mathbf{g}_{i}\|^{2}$.
+Mathematically, the diagonal sum represents the uncentered variance (update energy) of the samples, while the cross-terms correspond to the squared norm of the expected gradient $\|\mathbb{E}[\mathbf{g}_{\text{group}}]\|^{2}$.
+In RL for LLMs, the gradient noise variance inherently dwarfs the true gradient magnitude (i.e., $\text{Tr}(\text{Var}(\hat{A}_{i}\mathbf{g}_{i}))\gg\|\mathbb{E}[\hat{A}_{i}\mathbf{g}_{i}]\|^{2}$).
+Therefore, we can safely approximate the expected squared group gradient by its dominant diagonal components:
+$\mathbb{E}\left[\|\mathbf{g}_{\text{group}}\|^{2}\right]\approx\frac{1}{G^{2}}\mathbb{E}\left[\sum_{i=1}^{G}\hat{A}_{i}^{2}\|\mathbf{g}_{i}\|^{2}\right].$ (13)
+Let $V^{2}=\mathbb{E}[\|\mathbf{g}_{i}\|^{2}]$ denote the expected squared norm of the trajectory gradient (which corresponds to the trace of the Fisher Information Matrix).
+Under the standard simplifying assumption that the gradient magnitude is approximately independent of the specific binary correctness $r_{i}$, we can decouple the expectations.
+This elegantly isolates the advantage term, simplifying the expected squared group gradient to:
+$\mathbb{E}\left[\|\mathbf{g}_{\text{group}}\|^{2}\right]\approx\frac{V^{2}}{G^{2}}\mathbb{E}\left[\sum_{i=1}^{G}\hat{A}_{i}^{2}\right].$ (14)
+We now evaluate the advantage term $\sum_{i=1}^{G}\hat{A}_{i}^{2}$.
+In RLVR, verifiable rewards are binary, i.e., $r_{i}\in\{0,1\}$.
+Without standard-deviation normalization, the advantage is computed as $\hat{A}_{i}=r_{i}-u$, where $u$ is the expected group mean reward.
+The sum of squared advantages expands algebraically as:
+$\mathbb{E}\left[\sum_{i=1}^{G}\hat{A}_{i}^{2}\right]$ $=\sum_{i:r_{i}=1}(1-u)^{2}+\sum_{i:r_{i}=0}(-u)^{2}$
+$=c(1-u)^{2}+(G-c)u^{2}$
+$=Gu(1-u)^{2}+G(1-u)u^{2}$
+$=Gu(1-u)[(1-u)+u]=Gu(1-u).$ (15)
+Simultaneously, the Group Cumulative Advantage $\mathcal{A}_{\text{group}}$ is defined as the sum of absolute advantages:
+$\mathbb{E}\left[\mathcal{A}_{\text{group}}\right]$ $=\sum_{i=1}^{G}|\hat{A}_{i}|=\sum_{i:r_{i}=1}|1-u|+\sum_{i:r_{i}=0}|-u|$
+$=c(1-u)+(G-c)u$
+$=Gu(1-u)+G(1-u)u=2Gu(1-u).$ (16)
+Comparing Equation 15 and Equation 16 reveals a strict algebraic identity unique to the binary group-relative advantage formulation:
+$\mathbb{E}\left[\sum_{i=1}^{G}\hat{A}_{i}^{2}\right]\equiv\frac{1}{2}\mathbb{E}\left[\mathcal{A}_{\text{group}}\right].$ (17)
+Substituting Equation 17 into Equation 14, we obtain the final exact proportionality:
+$\mathbb{E}\left[\|\mathbf{g}_{\text{group}}\|^{2}\right]\approx\left(\frac{V^{2}}{2G^{2}}\right)\mathbb{E}[\mathcal{A}_{\text{group}}]=2Gu(1-u).$ (18)
+This demonstrates that the expected optimization step size (squared gradient norm) is strictly proportional to the Group Cumulative Advantage. ∎
+Theoretical Implications.
+Proposition B.2 fundamentally justifies our algorithmic design.
+By demonstrating that the expected variance (i.e., update energy) of the parameter gradient is mathematically locked to $\mathcal{A}_{\text{group}}$, we confirm that $\mathcal{A}_{\text{group}}$ acts as a rigorous surrogate for the gradient signal strength allocated to each problem.
+This mathematically verifies that existing GRPO variants intrinsically under-update hard problems (where $\mathcal{A}_{\text{group}}$ is distortedly low), directly motivating the necessity of our Difficulty Adaptive Rollout Sampling (DARS) methodology to re-balance update magnitudes.
+Appendix C Theoretical Connection Between DARS and Maximum Likelihood Reinforcement Learning
+Recent work on Maximum Likelihood Reinforcement Learning (MaxRL) (Tajwar et al., 2026) demonstrates that standard expected-reward objectives systematically under-weight high-difficulty instances.
+To counteract this, MaxRL introduces a modified advantage function to approximate the Maximum Likelihood (ML) objective.
+In this section, we provide a formal theoretical analysis demonstrating that our Difficulty-Adaptive Rollout Sampling (DARS) with the Hardness-Weighted (HW) schedule optimizes an expected objective identical to MaxRL.
+Crucially, we prove that while MaxRL relies on algebraic advantage scaling, which inherently exacerbates gradient variance.
+While DARS achieves the same objective through physical compute reallocation.
+We establish that DARS serves as a strict variance-reduced estimator of the ML-equivalent objective, offering superior optimization stability.
+C.1 Objective Equivalence via Cumulative Advantage
+We first analyze the weighting behavior of both algorithms utilizing our proposed Group Cumulative Advantage ($\mathcal{A}_{group}$) framework.
+Let $N$ denote the number of rollouts for a given prompt $q$, and let $u\in(0,1)$ denote the mean reward (i.e., empirical accuracy or pass rate), where the binary reward is $r_{i}\in\{0,1\}$.
+Proposition C.1 (Equivalence of Difficulty Weighting Profile) Both MaxRL and DARS-HW impose a problem-level optimization weight that is strictly proportional to (1u)(1-u).
+Consequently, both algorithms share an identical inductive bias: linearly up-weighting the optimization focus on problems as their difficulty increases ($u\to 0$).
+Proof.
+For MaxRL, the advantage is scaled as $\hat{A}^{MaxRL}_{i}=\frac{r_{i}-u}{u}$.
+For a group of $N$ rollouts, the expected number of successful trajectories ($r_{i}=1$) is $Nu$, and the expected number of failed trajectories ($r_{i}=0$) is N(1u)N(1-u).
+The cumulative advantage is thus computed as:
+$\mathcal{A}_{group}^{MaxRL}(u)$ $=\mathbb{E}\left[\sum_{i=1}^{N}\left|\hat{A}^{MaxRL}_{i}\right|\right]$
+$=Nu\cdot\left|\frac{1-u}{u}\right|+N(1-u)\cdot\left|\frac{0-u}{u}\right|$
+$=N(1-u)+N(1-u)=2N(1-u)$
+For DARS-HW, we employ the unscaled advantage $\hat{A}_{i}=r_{i}-u$. Based on our proposed schedule (Equation 7), the target cumulative advantage is explicitly constrained to:
+$\mathcal{A}_{group}^{HW}(u)=2(1-u)\mathcal{A}_{group}^{N_{pre}}(0.5)=C\cdot(1-u)$
+where $C=2\mathcal{A}_{group}^{N_{pre}}(0.5)$ is a constant independent of the prompt difficulty $u$.
+Since $\mathcal{A}_{group}^{MaxRL}(u)\propto(1-u)$ and $\mathcal{A}_{group}^{HW}(u)\propto(1-u)$, the weighting profiles are equivalent.
+∎
+C.2 Equivalence of Expected Gradients
+To formalize the alignment in optimization trajectories, we compare the total expected gradient contribution of a single prompt $q$ under both methods.
+Let $\nabla J_{i}(u)=(r_{i}-u)\nabla_{\theta}\log\pi_{\theta}(o_{i}|q)$ denote the standard baseline-subtracted policy gradient for a single rollout.
+Proposition C.2 (Equivalence of Expected Gradient Updates) Let $K$ be a positive scalar.
+The expected gradient updates of MaxRL and DARS-HW are proportional: $\mathbb{E}[\nabla J^{MaxRL}_{q}]\propto\mathbb{E}[\nabla J^{DARS-HW}_{q}]\propto\frac{1}{u}\mathbb{E}[\nabla J_{i}(u)]$.
+Both methods approximate the Maximum Likelihood gradient by universally scaling the standard policy gradient by a factor of $1/u$.
+Proof.
+For MaxRL, using a fixed rollout size $N$, the total expected gradient is scaled algebraically:
+$\mathbb{E}[\nabla J^{MaxRL}_{q}]=\mathbb{E}\left[\sum_{i=1}^{N}\frac{r_{i}-u}{u}\nabla_{\theta}\log\pi_{\theta}(o_{i}|q)\right]=\frac{N}{u}\cdot\mathbb{E}\left[\nabla J_{i}(u)\right]$
+For DARS-HW, we dynamically adjust the rollout size $N_{DARS}(u)$ to match the target $\mathcal{A}_{group}^{HW}(u)=C(1-u)$.
+Recall the standard group cumulative advantage property $\mathcal{A}_{group}=2N_{DARS}u(1-u)$.
+Equating the two yields $2N_{DARS}(u)u(1-u)=C(1-u)$, which simplifies to $N_{DARS}(u)=\frac{C}{2u}$.
+Thus, the expected gradient contribution is:
+$\mathbb{E}[\nabla J^{DARS-HW}_{q}]=\mathbb{E}\left[\sum_{i=1}^{N_{DARS}(u)}(r_{i}-u)\nabla_{\theta}\log\pi_{\theta}(o_{i}|q)\right]=\frac{C}{2u}\cdot\mathbb{E}\left[\nabla J_{i}(u)\right]$
+Up to the constants $N$ and $C/2$, both expected gradients scale strictly by $1/u$. ∎
+C.3 Variance Analysis: DARS-HW versus MaxRL
+While Proposition C.2 establishes that DARS-HW and MaxRL optimize the same expected objective, they diverge significantly in their construction of the Monte Carlo gradient estimator.
+Let $\Sigma_{g}=\text{Var}(\nabla J_{i}(u))$ denote the trace of the covariance matrix of the unscaled token-level gradient for a single rollout.
+We make the standard assumption that conditioned on the prompt $q$ and current policy $\pi_{\theta}$, individual rollouts are independent and identically distributed (i.i.d.).
+Let $u\to 0$ denote the regime of increasingly difficult reasoning tasks.
+For the MaxRL estimator, the variance relies on algebraic scalar multiplication over a fixed $N$ samples.
+By the properties of variance for i.i.d.
+variables ($\text{Var}(aX)=a^{2}\text{Var}(X)$):
+$\text{Var}(\hat{\nabla}J^{MaxRL}_{q})=\text{Var}\left(\sum_{i=1}^{N}\frac{1}{u}\nabla J_{i}(u)\right)=\sum_{i=1}^{N}\frac{1}{u^{2}}\text{Var}(\nabla J_{i}(u))=\frac{N}{u^{2}}\Sigma_{g}$
+Conversely, DARS-HW physically expands the sampling space, drawing $N_{DARS}(u)=\frac{C}{2u}$ independent trajectories while maintaining the bounded unscaled advantage $|r_{i}-u|\leq 1$:
+$\text{Var}(\hat{\nabla}J^{DARS-HW}_{q})=\text{Var}\left(\sum_{i=1}^{N_{DARS}(u)}\nabla J_{i}(u)\right)=\sum_{i=1}^{N_{DARS}(u)}\text{Var}(\nabla J_{i}(u))=N_{DARS}(u)\Sigma_{g}=\frac{C}{2u}\Sigma_{g}$
+Taking the ratio of the two variances yields:
+$\frac{\text{Var}(\hat{\nabla}J^{MaxRL}_{q})}{\text{Var}(\hat{\nabla}J^{DARS-HW}_{q})}=\frac{N/u^{2}}{C/(2u)}=\frac{2N}{C}\cdot\frac{1}{u}\propto\frac{1}{u}$
+Therefore, as $u\to 0$, the variance of MaxRL grows at a rate of $\mathcal{O}(1/u^{2})$, while DARS-HW grows at a significantly slower rate of $\mathcal{O}(1/u)$.
+C.4 Theoretical Analysis Conclusions
+  • Unified Optimization Objective with Divergent Implementation Routes.
+DARS-HW and MaxRL are two distinct implementation routes toward the same Maximum Likelihood reinforcement learning objective.
+MaxRL achieves the target optimization weight by algebraically modifying the advantage function ($\hat{A}^{MaxRL}_{i}=\frac{r_{i}-u}{u}$), scaling the gradient signal of each rollout by $1/u$.
+In contrast, DARS-HW retains the original unscaled advantage function ($\hat{A}_{i}=r_{i}-u$) and instead achieves the identical expected gradient scaling by adaptively increasing the number of rollouts ($N_{DARS}(u)\propto 1/u$) for difficult problems.
+These two approaches are mathematically equivalent in expectation but differ fundamentally in their implementation strategy.
+  • Variance Reduction of DARS.
+Appendix C.3 fundamentally explains the superiority of DARS over pure advantage-scaling methods as MaxRL.
+For complex reasoning tasks where the initial pass rate is extremely low ($u\ll 1$), MaxRL heavily amplifies the gradient of rare, potentially idiosyncratic correct trajectories.
+This leads to a high-variance gradient regime, which may inducing optimization instability and premature convergence.
+By contrast, DARS-HW functions as a variance-reduced estimator.
+Instead of amplifying a small fixed set of rollouts algebraically, DARS-HW allocates explicit computational budget to difficult problems, thereby expanding the breadth of actual correct reasoning paths discovered.
+This mechanism preserves a high signal-to-noise ratio in the gradient and maintains robust token-level entropy, theoretically validating DARS-HW as a more stable and efficient paradigm for Maximum Likelihood RL optimization.
 Appendix D Training and Evaluation Details